### 05 Feb 2024 v1

# Automatic Combination of Sample Selection Strategies for Few-Shot Learning

Abstract
In few-shot learning, such as meta-learning, few-shot fine-tuning or in-context learning, the limited number of samples used to train a model have a significant impact on the overall success. Although a large number of sample selection strategies exist, their impact on the performance of few-shot learning is not extensively known, as most of them have been so far evaluated in typical supervised settings only. In this paper, we thoroughly investigate the impact of 20 sample selection strategies on the performance of 5 few-shot learning approaches over 8 image and 6 text datasets. In addition, we propose a new method for automatic combination of sample selection strategies (ACSESS) that leverages the strengths and complementary information of the individual strategies. The experimental results show that our method consistently outperforms the individual selection strategies, as well as the recently proposed method for selecting support examples for in-context learning. We also show a strong modality, dataset and approach dependence for the majority of strategies as well as their dependence on the number of shots – demonstrating that the sample selection strategies play a significant role for lower number of shots, but regresses to random selection at higher number of shots.
[…]
1 Introduction
Many domains are characterised by a labelled data scarcity due to data collection and annotation costs and privacy, making the use of typical deep learning unfeasible. To address these problems, few-shot learning aims to train or adapt machine learning models to new tasks using only few labelled samples per class (Song et al., 2023). There are two main categories of few-shot learning (more details regarding their difference are included in Appendix B): 1) gradient-based, which includes meta-learning and fine-tuning, where the knowledge from relevant datasets is transferred to new problem using gradient-based training, similar to supervised learning (Song et al., 2023; Chen et al., 2019; Vanschoren, 2018; Hospedales et al., 2021); and 2) few-shot learning via in-context learning, a new paradigm, where a pretrained large language model is conditioned on few examples to generate the prediction without any parameter updates (Dong et al., 2022; Liu et al., 2023).
As the number of utilised labelled samples is small, the role of individual samples is all the more important. Based on the choice of samples for in-context learning, the observed performance can vary between state-of-the-art and random predictions, indicating a strong sensitivity to the sample selection (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022). Even though a lot of research is dedicated to sample selection, such as active learning (Ren et al., 2021) or core-set selection (Guo et al., 2022), the most common approach in few-shot learning is a simple random selection. The exception is in-context learning, where a large number of sample selection strategies have been proposed recently, selecting samples according to their properties such as similarity, diversity, informativeness or quality (Li & Qiu, 2023; Zhang et al., 2022; Chang & Jia, 2023).
However, the impact of sample selection strategies on in-context learning or gradient-based few-shot learning remains largely under-explored. Therefore, the goal in this paper is to investigate the impact of the sample selection strategies on few-shot image and text classification tasks across a wide range of datasets.
In addition, as the different strategies mostly focus on selecting samples based on a single property (e.g., how often the samples are forgotten (Toneva et al., 2018), how easy they are to learn (Swayamdipta et al., 2020; Zhang & Plank, 2021) or how informative they are (Ye et al., 2023b; Liu & Wang, 2023)), we ask whether the selection based on their combination would not lead to better results. To answer this question and to leverage the strengths of different sample selection strategies, we propose ACSESS, an effective method for Automatic Combination of SamplE Selection Strategies. The proposed method first identifies a subset of relevant strategies that can improve the overall success of few-shot learning. Afterwards, the identified strategies are combined together using weighting based on their expected contribution to find samples that can provide the most benefit. We summarise our contributions as follows111To support replicability and extension of our results, we openly publish the source code of our experiments at https://anonymous.4open.science/r/ACSESS:
We evaluate the impact of 20 sample selection strategies on 5 few-shot learning approaches, including meta-learning, few-shot fine-tuning and in-context learning, across 8 image and 6 text datasets. Experimental results show the selection strategies often lead to non-negligible improvements in performance (up to percentage points for gradient few-shot learning or percentage points for in-context learning), while showing a strong approach and dataset dependence.
We propose a method for automatic combination of sample selection strategies to leverage their strengths and identify samples with complementary properties required for the success of few-shot learning. Experimental results show the method consistently leads to performance improvement (up to percentage points for gradient few-shot learning or percentage points for in-context learning), outperforming all investigated strategies, including a recent in-context specific strategy LENS (Li & Qiu, 2023).
We further analyse how the impact of sample selection changes when increasing the number of shots. We find following key insights: 1) sample selection has significantly higher impact on lower number of shots; 2) at higher number of shots (30-40 on average), the impact of sample selection is negligible as all strategies regress to random selection; and 3) after a certain point, the boost in performance from increasing the number of shots is negligible (50-shot for gradient few-shot learning; 20-shot for in-context learning with strong dependence on input text size and model context size).
2 Related Work
A lot of focus is dedicated to sample selection: 1) to reduce the computation costs without the loss of performance, by distilling dataset to lower number of samples (Yu et al., 2023); 2) selecting a core set of a dataset with similar properties and same performance as the full set (Guo et al., 2022); or 3) to reduce annotation costs using active learning (Ren et al., 2021). A specific research focus is dedicated to selecting good performing samples for in-context learning, as the overall performance was found to be sensitive to this choice, leading to large variability in results (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022). However, the adoption of sample selection in few-shot learning is limited to active learning strategies and in-context learning.
The sample selection strategies can be divided into two groups. The first, more prevalent, uses heuristics and unsupervised metrics to rate and select samples according to specific property that serves as estimate of their potential to increase the model’s performance. The most popular approach is selecting samples based on the similarity to the test sample for in-context learning. The methods differ in employed: 1) similarity measure, such as cosine/Euclidean distance (Liu et al., 2022; An et al., 2023; Zemlyanskiy et al., 2022; Gupta et al., 2022; Pasupat et al., 2021), BM-25 (Wang et al., 2022; Nashid et al., 2023) or their combination (Agrawal et al., 2023; Gao et al., 2021); and 2) representation, such as simple TF-IDF (Agrawal et al., 2023; Wang et al., 2022; Zemlyanskiy et al., 2022), word embeddings from basic or further fine-tuned transformers (Liu et al., 2022; Gao et al., 2021; Gupta et al., 2022; Pasupat et al., 2021), or representation generated by prompting a large language model (An et al., 2023; Nashid et al., 2023).
Besides similarity, the samples are often selected based on their informativeness or the uncertainty. The active learning strategies are most common (Köksal et al., 2022; Schröder et al., 2022; Margatina et al., 2021; Park et al., 2022; Mavromatis et al., 2023), but also other approaches are utilised, such as using the bias of the samples (Ma et al., 2023). Other methods define a notion of quality for each sample, either by prompting a large language model to rate the samples (Shin et al., 2021) or observing how the inclusion or removal of the sample affects the performance (Maronikolakis et al., 2023; Nguyen & Wong, 2023).
When considering representativeness, the samples are selected using core-set methods that are designed to find a subset of samples with properties that are representative of the full training set (Guo et al., 2022; Killamsetty et al., 2021; Mirzasoleiman et al., 2020; Paul et al., 2021; Coleman et al., 2020). Learnability metric either determines how easy it is to learn the specific sample (Swayamdipta et al., 2020; Zhang & Plank, 2021), or how often the sample is forgotten after being learned (Toneva et al., 2018).
Finally, some approaches balance multiple properties at the same time. Often, the similarity of samples is balanced with their diversity (Wu et al., 2023; Qin et al., 2023) or the informativeness of the samples is balanced with their representativeness (Su et al., 2022; Levy et al., 2023; Ye et al., 2023b; Liu & Wang, 2023). In specific cases, the search is done in two step fashion, such as finding set of informative samples and then using diversity guided search to improve this set (Li & Qiu, 2023).
The second set of approaches leverages training to select the samples. One possibility is to train a sample retriever, such as a separate scoring model (Rubin et al., 2022; Luo et al., 2023; Li et al., 2023; Wang et al., 2023; Aimen et al., 2023) or using determinantal point process (Yang et al., 2023; Ye et al., 2023a). Another set of approaches leverages reinforcement learning to select good samples (Zhang et al., 2022; Shum et al., 2023; Scarlatos & Lan, 2023). Finally, the Datamodels approach trains a linear model to predict the performance gain of a set of samples to select the subset that would lead to the highest possible performance increase (Ilyas et al., 2022; Chang & Jia, 2023; Jundi & Lapesa, 2022; Vilar et al., 2023).
In this work, we investigate the impact of the existing strategies on few-shot setting. In contrast to (Li & Qiu, 2023), which represents a work closest to ours, we explore multiple few-shot learning approaches not only in-context learning. Our proposed method is inspired by the Datamodels approach (Ilyas et al., 2022), but performs the selection at the level of strategies instead of samples.
3 Sample Selection Strategies
3.1 Single-Property Strategies
| Strategy | Description | |
Informativeness |
Similarity |
Most similar samples (popular for in-context learning). |
Diversity |
Most diverse/dissimilar samples. |
Entropy (Park et al., 2022) |
Active Learning; Highest entropy over class probabilities. |
Margin (Park et al., 2022) |
Active Learning; Lowest difference in probability between top 2 classes. |
Least Confidence (Park et al., 2022) |
Active Learning; Top class with lowest probability assigned. |
Loss (Park et al., 2022) |
Active Learning; Highest loss. |
Contrastive Active Learning (CAL) (Margatina et al., 2021) |
Core-Set; Predictive likelihood that diverges most from the neighbourhood as determined by KL divergence. |
GraNd (Paul et al., 2021) |
Core-Set; Highest contribution to decline of training loss. |
Core-Set; Submodularity function that measures diversity and informativeness. |
Representativeness |
Core-Set; Minimising distance of subset centre to full dataset. |
Core-Set; Minimising distance of every sample in subset to full dataset. |
CRAIG (Mirzasoleiman et al., 2020) |
Core-Set; Gradients representative of full dataset. |
Glister (Killamsetty et al., 2021) |
Core-Set; Bi-level optimisation. |
Learnability |
Forgetting (Toneva et al., 2018) |
Core-Set; How often the samples are forgotten. |
How easy it is to learn the samples. Due to no obvious consensus, all of easy, ambiguous, hard and combination of easy and ambiguous samples are considered. |
In this work, we thoroughly investigate the impact of strategies that consider 3 groups of properties: 1) informativeness, or how informative the samples are for the model; 2) representativeness, or how well the subset of samples represents the full dataset; and 3) learnability, or how easy it is to obtain or retain the information contained in the samples. We specifically focus on strategies that perform the selection by assigning a score to each sample that is used to sort the samples and choose the best. This includes active learning strategies that measure the uncertainty of the model, core-set selection strategies that find subset representative of the full-set and other strategies that use training dynamics. The high-level overview of these strategies is presented in Table 1 and their further details are included in Appendix C.
3.2 ACSESS: Automatic Combination of Strategies
Even though the different selection strategies can identify samples with specific single property, the use of these identified samples may not have a significant impact on the performance when working in a few-shot setting. For example, an informative sample that is a hard to learn or is often forgotten, may not contribute as much, leading to low impact on the performance. Instead, we propose to select examples that are characterised by a balance of complementary properties, even when they do not represent the best sample of any given property (e.g., a sample that may be less informative, but is more easily learned or retained). To accomplish this, we propose ACSESS, an effective method for Automatic Combination of SamplE Selection Strategies, composed of two stages: 1) identification of a subset of well-performing and relevant strategies that can improve the overall success of few-shot learning; and then 2) combination of the identified strategies using a weighted scheme based on their expected contribution. The final set of samples is selected by sorting the samples according to their final score and selecting the ones with the highest score.
Identifying the Subset of Relevant Strategies
In the first phase, the aim is to identify relevant strategies that, when combined, can select high quality samples based on their different properties. Even though the most optimal combination would be achieved by searching through all the combinations of strategies, such search is infeasible as it requires extensive computation costs. Instead, we use three separate methods for selecting strategies, inspired by feature selection in traditional machine learning and the Datamodels approach (Ilyas et al., 2022), specifically: 1) forward strategy selection; 2) backward strategy selection; and 3) datamodels strategy selection. The final subset of selection strategies is composed of the ones that are identified by each of the methods as most impactful, to minimise the effect of potential drawbacks of the individual method.
Forward Selection iteratively adds to the subset the selection strategy that results in the highest performance increase (similarly to forward feature selection). Starting with an empty subset, the method alternates between two steps. First, for each strategy not included in the subset we determine the increase in performance the strategy would yield if added to the subset (evaluating the performance of samples selected by the combination of already selected subset and the potential new strategy and comparing to old subset). Afterwards, the strategy that provides the highest positive increase in performance is added to the subset and the first step is repeated. If no added strategy leads to increase of the overall performance, the process ends, resulting in the subset of relevant selection strategies.
Backward Selection iteratively removes strategies from the subset until the performance no longer increases (similarly to backward feature selection). The method works the same way as the forward selection, but starting with all selection strategies and removing them from the subset based on the performance increase such removal leads to.
Datamodels Selection draws inspiration from the Datamodels (Ilyas et al., 2022; Chang & Jia, 2023), but works on the level of the strategies instead of samples. First, a set of 150 random combinations of strategies is created and evaluated. The random search is constrained to include each strategy at least 5 times to ensure sufficient coverage of each strategy. Afterwards, these combinations are used to create a simple regression dataset, with the difference in performance to the baseline selection (the classic few-shot learning selection) as label, and the multi-label presence vector of strategies as features. The dataset is used to train a regularised linear regression, LASSO (Tibshirani, 1996), which zeros out the weights for as many strategies as possible. The selection strategies with positive weights in the model represent the subset of relevant strategies.
Combining the Identified Strategies
After the relevant selection strategies are identified, they are combined using a weighting scheme. Each sample is assigned an importance score by each of the identified strategies. As the different strategies may assign scores of different magnitudes, all scores are normalised to zero-one interval. The final sample score is obtained as a weighted average of its assigned scores and the samples with highest score is selected.
In this paper, we evaluate three separate weighting schemes. First, uniform combination, assigns the same weight to each of the identified strategies. This represents the simplified version of the ACSESS method, potentially resulting in lower performance impact, but also with lower computation costs as no further optimisation is required. This weighting is also used for forward and backward selection. Second, weighted combination, assigns a different weight to each strategy, prioritising specific sample properties. The final trained weights of the Datamodels linear model are used as the dataset and model specific weights for the different strategies. As the weighting scheme is dataset and model specific, it should lead to the highest performance impact, at the cost of requiring more extensive computation. Finally, the weighted combination with random selection introduces a random element to the weighted combination, as random selection was identified as a strong baseline that can improve active learning performance (Park et al., 2022). Each sample is assigned additional uniformly randomised score which is included in the combination. The weight of this score is determined by the Datamodels by including random selection as additional strategy, due to its impact being dataset and model dependent.
4 Experiments
Datasets. The experiments are conducted on image and text classification datasets composed of different tasks with different number of classes. For the image datasets, we use 8 randomly chosen datasets from the Meta-Album (Ullah et al., 2022), specifically LR_AM.DOG and LR_AM.AWA for animal classification, VCL.APL for plane classification, HUM_ACT.ACT_410 for classification of human actions, MNF.TEX_DTD for classification of texture, MCR.PRT and MCR.PNU for classification of microscope images, and PLT.FLW for flower classification. In addition, we use 6 widely-adopted text datasets: 20 News Group (Lang, 1995) and News Category (Misra, 2022) for news category classification and ATIS (Hemphill et al., 1990), Facebook (Schuster et al., 2019), HWU-64 (Liu et al., 2021) and SNIPS (Coucke et al., 2018) for intent classification.
Few-shot Learning Approaches. The evaluation is done in a 5-way 5-shot setting using the gradient few-shot learning on both the image and text datasets and the in-context learning on the text datasets. For gradient few-shot learning, we use meta-learning approaches, specifically Prototypical Networks (Snell et al., 2017) and MAML (Finn et al., 2017), and Few-Shot Fine-Tuning (Chen et al., 2019). For in-context learning, we use the Mistral-7B (Jiang et al., 2023) and Zephyr-7B (Tunstall et al., 2023) models.
Baselines. The selection strategies are compared with 2 baselines: 1) Classic selection, where a new set of samples is randomly selected for each new task, potentially covering the whole datasets when evaluating on enough tasks; and 2) Random selection, where only a single set of 5 samples per class is randomly selected at the start and used for every single single task, representing a more realistic setting with limited budget. In-context learning is also compared with the recently proposed LENS (Li & Qiu, 2023) strategy that is in-context learning specific. Each selection strategy is run over 10 random seeds to determine their sensitivity to the effects of randomness (e.g., randomly initialised seeds). In addition, each experiment is repeated 5 times for different data splits and an average accuracy over these runs is reported (if not specified otherwise). Further experimental setup details are in Appendix D.
4.1 Impact of Single-Property Selection Strategies
| Strategy | Image Data - FSL | Text Data - FSL | Text Data - ICL | |||||
| ProtoNet | MAML | FT | ProtoNet | MAML | FT | Mistral | Zephyr | |
| Random | ||||||||
| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||
| Similarity | ||||||||
| Diversity | ||||||||
| Entropy | ||||||||
| Margin | ||||||||
| Least Confidence | ||||||||
| Loss | ||||||||
| CAL | ||||||||
| CRAIG | ||||||||
| DeepFool | ||||||||
| Forgetting | ||||||||
| Glister | ||||||||
| GraNd | ||||||||
| Herding | ||||||||
| KCenter | ||||||||
| Graph-Cut | ||||||||
| CartographyEasy | ||||||||
| CartographyAmbiguous | ||||||||
| CartographyHard | ||||||||
| CartographyEasy+Ambig. | ||||||||
| ACSESSUniform | ||||||||
| ACSESSWeighted | ||||||||
| ACSESSWith Random | ||||||||
In this section, our goal is to answer the following research question: RQ1: What is the impact of single-property selection strategies on the few-shot learning? The objective is to analyse the performance increase of the different selection strategies across various datasets and few-shot learning approaches. The results are shown in Figure 1 and Table 2, aggregated over all the datasets (the detailed de-aggregated results are included in Appendix F). We present the results separately for image and text datasets, as well as for the gradient few-shot learning and in-context learning, as we observed a strong dependence of the strategy benefit on the modality and few-shot learning approach used.
Majority of selection strategies fail to outperform the Classic selection in case of gradient few-shot learning. In most cases, the strategies lead to performance on par or even lower than the Classic selection (for example hard samples on image datasets or diverse samples on text datasets). Only specific strategies consistently lead to better performance (e.g., easy to learn samples on image datasets, or Glister on text datasets). As such, the random sample selection represents a strong baseline for the gradient few-shot learning.
Sample selection is more beneficial for in-context learning. Larger number of strategies leads to increase in performance, while only few strategies, such as the popular selection based on similarity or diversity, under-perform the Classic selection. At the same time, the increase in performance is higher than with gradient few-shot learning, although with a larger variance in results. In contrast to gradient few-shot learning, the hard to learn samples provide the most benefit for in-context learning. The reason may be that hard to learn samples provide a lot of information beneficial for the classification, which cannot be easily leveraged in the gradient few-shot learning due to the samples being hard to learn. However, as there is no explicit “learning” in the in-context learning use of large language models, such samples can be more easily leveraged.
Learnability property is a strong indicator of quality for few-shot learning. The strategies that choose based on how easy it is to learn the samples (Cartography) or how often they are forgotten (Forgetting) consistently lead to overall highest increase in performance. The observed improvement is on average as high as percentage point for gradient few-shot learning and percentage points for in-context learning.
4.2 ACSESS Results and Characteristics
In this section, we answer the following research question: RQ2: How does the ACSESS method perform in comparison to the single-property strategies? Besides analysing the benefit and comparing with other strategies, we analyse the influence of the identified strategies and of the different weighting schemes.
The ACSESS method consistently leads to performance increase across all few-shot learning approaches. Our proposed method outperforms all the investigated single-property strategies, leading to an average performance increase of up to percentage points for gradient few-shot learning and percentage points for in-context learning. Comparing with the in-context specific strategy LENS (Li & Qiu, 2023) (see Figure 2 for comparison), which also outperforms all single-property strategies, we observe the increase of ACSESS is statistically more significant (p-value of using the Wilcoxon signed-rank test).
The ACSESS method identifies samples with complementary properties, with higher priority on learnability. The identified selection strategies that are used in the combination are different for the different data modalities (image vs. text) and the type of few-shot learning (gradient vs. in-context learning). For the image datasets, the combination of easy to learn samples (Cartography) that are located on the decision boundary (Margin) provides the most benefit. For the gradient few-shot learning approaches on text datasets, the combination of easy and ambiguous samples (Cartography) that are least often forgotten (Forgetting) on the decision boundary (Margin) that provide additional informativeness from the Graph-Cut strategy provides the most benefit. In case of in-context learning, the combination of hard to learn samples (Cartography) that are least often forgotten (Forgetting) on the decision boundary (Margin) and with highest Entropy provide the most benefit. Further analysing the identified strategies and their weights, we find the ACSESS method gives higher importance to samples that are learnable, lower importance to informative samples (but with multiple sources for both), while the strategies that focus on representativeness are not included at all.
The dataset and approach specific weighting leads to the best performance. The only exception is the Few-Shot Fine-Tuning approach, where the uniform weighting shows the most benefit. At the same time, the weighting strategy that incorporates the random selection often leads to worse performance and higher variance in the results. Even though the uniform combination is computationally less expensive than the weighted combination, its performance increase is only slightly lower (average difference of ). As such, the uniform weighting represents a good trade-off between the performance increase and the computation costs, making it in general a better choice.
4.3 Effects of Variable Number of Shots
In this section, we answer the following research question: RQ3: How does the benefit of sample selection change based on the number of shots? Our aim is to investigate how the benefit of sample selection strategies changes as we increase or decrease the number of shots that we select. In addition, we investigate whether the few-shot learning approaches themselves benefit from larger number of shots. The results of this evaluation are included in Figure 3, comparing the benefit of the ACSESS strategy and Random selection over the main experimental setup (5-shot setting).
The selection strategies provide more benefit on lower number of shots. On the 1-shot setting, the difference in accuracy between the Random strategy and the ACSESS strategy can be as high as , or percentage points respectively for the few-shot learning on image data and text data, and in-context learning on text data. Increasing the number of selected shots leads to diminished impact, as the ACSESS strategy regresses to random selection when selecting more than 30-40 shots. For more detailed analysis we compare all the investigated strategies in the 5-shot and 10-shot setting (distribution of results is shown in Figure 5 in Appendix E). This comparison further reinforces our findings, as we find that the difference to random selection is statistically significantly higher in the 5-shot (p-value of using the Wilcoxon signed-rank test).
Impact of selection strategies and further increase of the number of shots is negligible after certain point. For the gradient few-shot learning, the performance after approximately 50 shots remains more or less constant. For the in-context learning, the performance even decreases significantly after 20-25 shots, achieving similar performance to the 1-shot setting when using 50-shots. The main reason for this drop-off may be the limited context size of the large language models, making the model degenerate to random guessing as the cross-over is text length dependent. As such, the focus on selecting high quality samples is paramount when only few labelled samples are available (e.g., working with limited labelling budget), while the random selection can be safely used with enough labels. Furthermore, labelling more samples and using them as additional shots for few-shot learning may not lead to better performance.
4.4 Additional Insights to Sample Selection Strategies
In this section, we provide additional insights from ablation studies. We are mainly interested how the factors such as dataset and model affect the sample selection (based on dataset specific results from Appendix F) and how sensitive the strategies are to their random components (based on the variance from repeated strategy runs in Figure 4).
Specific strategies show sensitivity to the effects of random seed. Changing the random seed for specific strategies may often lead to different set of selected samples and different performance impact. The only exception are some of the core set selection strategies (such as Glister or the Graph-Cut), which consistently lead to the selection of the same set of samples. On the other hand, strategies with random initial condition, such as the similarity selection or some of the active learning strategies, are more sensitive to the effects of the random seed. The ACSESS method shows lower sensitivity to repeated runs (with the exception of weighting with random element) but still may lead to different sets of samples based on the random seed.
Strong dataset dependence for selection strategies. The strategy that performs well on one dataset (such as DeepFool on the HUM_ACT.ACT_410 dataset) may perform quite poorly on other datasets (such as MCR.PRT or MCR.PNU). Such strategies cannot be used on any dataset without further evaluation. Only few strategies perform consistently on the majority of the datasets, such as Herding showing poor performance on all image datasets, or Cartography showing good performance across all datasets and models. We do not observe a significant dataset dependence of the ACSESS method, as it identifies the well-performing strategies and leverages their strengths for the specific dataset.
Strong model dependence and sensitivity to selected samples. The strategy performing well for one model (GraNd on MCR.PNU for MAML, or KCenter on 20 News for Zephyr) may result in a decrease of performance for other models (GraNd for Few-Shot Fine-Tuning, or KCenter for Mistral on the same setting). Overall, the selected samples resulting from different strategy or from a repeated run of the same strategy affect the Few-Shot Fine-Tuning less, leading to smaller benefit. At the same time, the MAML method and the in-context learning models are more affected by the selected samples, leading to more varied results and often higher performance increase. Such sensitivity to sample selection was observed in previous works as well (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022; Li & Qiu, 2023). Curiously, Prototypical Networks gain quite large performance increase from sample selection, but show low sensitivity to the selected samples and variance from repeated runs of sample selection. As such, sample selection is more important for MAML and In-Context Learning and less important for Few-Shot Fine-Tuning.
5 Conclusion
In this paper, we investigate the impact of larger number of different sample selection strategies on the success of few-shot learning. To leverage strengths of these strategies, we propose the ACSESS method for automatic combination of the sample selection strategies to select samples with complementary properties. The experimental results show that although the individual sample selection strategies may not always be beneficial, their combination using our proposed method consistently leads to better performance. Curiously, we observe that the learnability of the samples is more important for few-shot learning than their informativeness or representativeness. Finally, analysing the behaviour of the sample selection in regards to the number of selected shots, we find that the sample selection provides the most benefit on low number of shots, while at the higher number of shots all strategies regress to random guessing. In addition, after certain point, increasing the number of available samples no longer leads to improvement in performance of few-shot learning, but may lead to decrease in performance in some cases, especially for in-context learning.
Acknowledgements
This work was partially supported by the projects funded by the European Union under the EU Horizon 2020: TAILOR, GA No. 952215, by the European Union under the Horizon Europe: vera.ai, GA No. 101070093, AI-CODE, GA No. 101135437; and by Modermed, a project funded by the Slovak Research and Development Agency, GA No. APVV-22-0414.
Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of Society, Operational Program Integrated Infrastructure.
Ethical Considerations and Impact Statement
The experiments in this paper work with publicly available datasets, citing the original authors of each dataset. All the datasets are used in accordance to their license and terms and conditions of use. We do not work with any personally identifiable information or offensive content and perform no crowdsourcing or further data annotation. In addition, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the advancement of the field of Machine Learning and Few-Shot learning. Finally, we follow the license terms for all the models we use – all models and datasets allow their use as part of research. It is possible the large language models (Mistral and Zephyr) we used contain biases and may generate potentially offensive or harmful content. However, the authors of the models already reduce this potential bias as much as possible when training the models, while at the same time we limit the output to few tokens and do not release any output of these models, which should further reduce the potential bias and negative impact.
Impact Statement: CO2 Emissions Related to Experiments
The experiments presented in this paper used higher compute resources in order to evaluate all the sample selection strategies for all the different few-shot learning approaches. Overall, the experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCOeq/kWh. A cumulative of 2000 hours of computation was performed on hardware of type A100 PCIe 40GB (TDP of 250W). Total emissions are estimated to be 216 kgCOeq of which 0 percent were directly offset. These estimations were conducted using the MachineLearning Impact calculator presented in (Lacoste et al., 2019). Whenever possible, we tried to reduce the compute resources used as much as possible. The most compute resources were used by the large language models – Mistral-7B and Zephyr-7B and their further tuning. To reduce the computation costs, we use the 4-bit quantised versions of these models and in cases when the further tuning is required, we fine-tune only the last classification layer (keeping the rest frozen). In addition, for gradient few-shot learning approaches, we opt for smaller models that do not require such an extensive computation costs, while still providing relevant results. Finally, whenever possible, we reuse the already performed computations for the selection strategies, such as generating the feature representation for the different samples only a single time and re-using it whenever possible (e.g., using pre-trained models without further tuning for these representations).
Appendix A Limitations
In the experiments for evaluating the impact of different sample selection strategies, we focused on such settings, where the impact of sample selection may be reduced — despite that we observed significant differences in the achieved performance. First of all, in within-domain few-shot learning, the samples are selected from the same domain, thus the difference in impact and role of the individual samples may not be as high. We may expect that the impact of sample selection in cross-domain few-shot learning may be higher and lead to even larger performance deviations. In addition, we focus on a balanced-shot setting, where we select the same number of shots for each class, leading to lower impact of strategies for different classes. These additional settings are out of scope of this paper and may be a subject of future research.
To diminish undesirable randomness effects, we report the results as an average over 5 repeats of data splitting, 10 repeated runs of the selection strategies (resulting in 50 overall evaluation runs), with each run evaluated on 600 tasks. Such setting limits the variance in results, but may also obscure the benefit of selection strategies to a certain extent. On the other hand, although we focus only on a 5-way few-shot learning setting (using only 5 classes per task), we do not expect significant change in result for most of the few-shot learning approaches when using higher number of classes (such as 10-way or 20-way classification). The only exception is the in-context learning, where we observed the dependence on input size and context width of the models. With higher number of classes, the observed performance of large language models would start to deteriorate to the performance that can be achieved with a low number of shots.
Similarly to other research papers focusing on sample selection and few-shot learning via in-context learning, we assume an availability of large-enough annotated dataset from which we select the sample using different strategies. This also includes the use of larger validation or held-out set when selecting strategies. If we considered the true few-shot learning setting, only the active learning strategies (i.e., Entropy, Margin and Least Confidence) could be used, as they work with unlabelled data only, or a different approach for sample selection should be employed, such as using weak-labels to select the best samples, or assign every single class to all the labels and perform the sample selection that way to select both best performing samples as well as the class for them (as it is done in (Chang & Jia, 2023)).
Before the experiments, we run hyperparameter optimisation for the gradient few-shot learning models and perform basic prompt-engineering for the in-context learning approaches. We follow the recommendations for good prompt engineering, using the meta-tags defined for each model (which lead to better performance of the models), while also not using overly-simple or overly-complicated prompts. The prompts we use are created based on dataset description, the prompts used in related works and the formats recommended for different models (e.g., taking inspiration from (Sun et al., 2023)). As we do not perform any extensive or automatic prompt-engineering, these prompts may lead to lowered performance and in some cases also lower impact of the sample selection for the few-shot learning setting. However, running an extensive prompt-engineering would introduce another controllable parameter which could potentially introduce further bias into the results (e.g., different selection strategies showing different impact for different prompt formats).
Finally, we also identify factors that we consider to be sufficiently covered, but where the further extension could lead to more generalisable results (nevertheless, also requiring a considerable amount of additional computational resources), such as including more selection strategies, evaluating on more datasets covering other classification tasks (mainly in regards to text datasets) or using larger number of base models for meta-learning and few-shot fine-tuning, or large language models for in-context learning. First of all, although we focus on the major selection strategies, which are general for any few-shot learning approach, the work can be extended to include larger number of in-context learning specific strategies or strategies that do not perform selection based on scoring of the samples. In addition, even though we run the experiments on 8 image datasets and 6 text datasets, covering different classification tasks with different number of classes, the work could be potentially extended to even larger number of datasets and classification tasks, mainly in case of text modality where we are limited by the number of classes the datasets contain (to sufficiently explore few-shot learning). Finally, although we cover all the main groups of few-shot learning approaches, the work could be extended to cover larger number of meta-learning approaches (where we focus on two of them), base models for meta-learning and few-shot fine-tuning (we focus on one model for image data and one model for text data) and large language models for in-context learning (we focus on two of the most recent and currently best performing large language models Mistral and Zephyr). However, we do not expect significantly different results and findings, only an improvement of their generalisability.
Appendix B Few-Shot Learning Problem Formulation
In this section, we introduce the definition of the few-shot learning problem and the difference between gradient-based few-shot learning and few-shot learning via in-context learning. Note that in-context learning can be applied only to text datasets as it utilises pretrained large language models.
For the standard few-shot learning, we assume a larger collection of instances , with larger number of classes. From this dataset, a separate task is created by randomly sampling few classes . The task is characterised by a training set (also called support set, ) and test set (also called query set, ). To create these sets, few samples () are sampled for each class. The few-shot learning algorithms then use the few samples from the support set to learn to solve the task quickly and are evaluated on this ability using the query set . This represents the -way -shot classification. The few-shot learning approaches mostly consist of two phases, training phase and testing phase. Often, these two phases work with separate sets of , i.e., the classes appearing as part of tasks in the testing phase do not appear during training. As each task works only with subset of classes, the evaluation is often done on many such tasks and the average performance from these tasks is reported.
The gradient-based few-shot learning approaches follow episodic training similar to typical supervised training to gather knowledge and experience and learn to quickly adapt when encountering new task. Using the support set, the model is adapted to the task using gradient descent based approaches and then evaluated on the query set. Based on the evaluation, some kind of knowledge on how to better adapt is saved in part of the model, while the rest of the model is reset. This is done for a large number of episodes (often called meta-training), with the model accumulating the knowledge. After enough episodes are run, the model is evaluated (in a meta-testing phase) using the exact same setup.
On the other hand, no training phase is present in few-shot learning via in-context learning, but a pretrained large language model is used without any additional parameter updates instead. For each new task, a task-specific text prompt is formed, by concatenating task-specific instructions, the samples from the support set along with the their labels, and a specific test sample. After being given this prompt, the large language model generate a prediction in a form of word (or multiple words). These generated words are then mapped to the task labels and a class is assigned. The process is repeated for each sample in the query set. For example in case of binary classification, the task-specific instruction may be ”Determine sentiment of following sentence”, the concatenated samples from support set may be formatted as ”sentence, sentiment: positive/negative”, with the word ”positive” mapping to class representing positive sentiment and ”negative” to class representing negative sentiment. As such, the only supervision for the large language model is from the instruction and the concatenation of support samples. Therefore, the model is limited by its max input length. At the same time, this approach was found to be sensitive to the selected samples in the support set and their order (Zhao et al., 2021; Zhang et al., 2022; Pecher et al., 2023; Chang & Jia, 2023).
Appendix C Sample Selection Strategies Details
In this section, we provide further details about the different single-property sample selection strategies that we evaluate in this paper, as well as for the baselines. In addition, we specify the hyperparameters of these methods and how we use them for selecting the set of samples.
The idea behind the uncertainty-based and active learning selection strategies is that the most informative samples with higher impact are the ones that the model is least certain about (Coleman et al., 2020; Park et al., 2022). These strategies select samples iteratively over multiple steps. In our experiments, we select 1 sample for every class in individual step. In addition, the first sample for every class is chosen randomly, as such setting was observed to lead to better active learning results (Park et al., 2022). When using these methods for combination, the samples that are selected after running all the required steps (e.g., 4 steps when choosing 5-shots) use the score assigned to them at their respective step when they were chosen, while the remaining samples are assigned the average score over all the selection steps. The Entropy strategy measures the entropy for each of the samples over the predicted class probabilities (which is used as the sample score) and selects the sample with highest entropy. The Margin (Breaking Ties) strategy assigns to each sample the difference between the probability of the two most probable classes. The samples with lowest such difference are selected as those are the ones the model is most uncertain about (and which can be considered on the decision boundary of the model). The Least Confidence strategy simply selects the samples where the most likely class has the lowest probability assigned to it. Finally, the Loss strategy selects samples with the highest loss.
The idea behind the core-set selection strategies is to select subset samples that are representative of the whole dataset and have similar properties. The included strategies select samples based on different criteria, such as error/loss, gradient or submodularity. These strategies select the full set of samples at the same time according to the assigned scores. As such, when using these strategies for combination, the full sample scores are used from this single iteration. The Decision Boundary based methods select samples the are close to the decision boundary, as these represent the samples that are hard to separate and should provide the most information to the models. The Contrastive Active Learning (CAL) (Margatina et al., 2021) selects the samples whose predictive likelihood diverges the most from their neighbourhood as determined by the KL divergence. The DeepFool strategy (Ducoffe & Precioso, 2018) approximates the samples on boundary by perturbing the samples until the predicted class changes and selects the samples that require the smallest perturbation. The idea behind Error or Loss based methods is that the samples that contribute more loss or error to the training of model have higher importance and should contribute more to the training when selected. The GraNd strategy (Paul et al., 2021) calculates the average contribution of each sample to the decline of the training loss at early stage of training across several independent training runs (calculating gradient norm expectation of the sample). The Forgetting strategy (Toneva et al., 2018) simply calculates how often the specific sample is incorrectly classified after being classified correctly in the previous epoch. The number of such forgetting events is used as the sample score and samples that are forgotten the most often (or least often) are selected. In our evaluation of single-property strategies, we consider only the setting of most often forgotten samples. However, when using the strategy for combination, we also consider the setting of selecting the least forgotten ones. The Submodularity based methods use submodular functions (Iyer & Bilmes, 2013; Iyer et al., 2021), which naturally measure the informativeness and diversity of a subset of samples. We evaluate the selection based on Graph-Cut, as the method performed the best from all the different submodular functions. The Geometry based methods assumes that samples that are close to the feature space have similar properties. As such, these methods try to remove samples that are close to each to as they provide redundant information and lead to small subset that covers the properties of the whole dataset. The Herding strategy (Welling, 2009; Chen et al., 2010) iteratively and greedily adds one sample to the subset which minimises the distance between the centre of the subset and the centre of the whole dataset. On the other hand, the KCenter strategy (Sener & Savarese, 2018; Agarwal et al., 2020) selects sample that minimises the largest distance between each sample already in the subset and its closest sample not yet included in subset. The Gradient Matching based methods try to find a subset of samples whose gradient can imitate the whole dataset. The CRAIG strategy (Mirzasoleiman et al., 2020) converts the problem into maximisation of a monotone submodular function which is then solved greedily. Finally, the Bilevel Optimisation based methods convert the problem into a bilevel optimisation, with different objectives in the outer and inner level. The Glister strategy (Killamsetty et al., 2021) uses a validation set on the outer optimisation, where the subset selection is used as objective, and log-likelihood in the inner optimisation, which handles model parameter optimisation.
In addition, we also evaluate and use sample selection based on the Cartography (Swayamdipta et al., 2020; Zhang & Plank, 2021), which measures how easy or hard it is to learn different samples. This ease of learning is determined by observing the training dynamics of all the samples across few epochs, looking at the average confidence/probability of the correct class and the variance of this confidence. The samples with high confidence and low variance are considered to be the easy to learn samples. At the same time the samples with small confidence and small to medium variance are considered the hard to learn ones. The remaining samples are considered to be ambiguous (medium confidence or samples with high variance). We explore four different settings in our experiments and choose the samples accordingly: 1) Easy samples, where we sort the samples based on confidence and choose the top-K samples with highest confidence for each class; 2) Hard samples, where we sort the samples based on confidence and choose the bottom-K samples for each class (i.e., the lowest confidence samples); 3) Ambiguous, where we first calculate average confidence and standard deviation of all the samples, select the samples whose confidence is close around the average confidence (in the interval defined by 1 standard deviation around the average confidence), and then randomly sample from them; and 4) Easy + Ambiguous, where we choose half of the samples from the easy set and the other half from ambiguous set.
All the approaches mentioned so far perform the selection using a model that is trained for some epochs. The number of epochs the training is done for significantly affects the outcome of different strategies. For this reason, we search for an optimal number of epochs by evaluating the selected sample on a single run. The optimal number of epochs determined by this search is 10% of the full epochs used for the few-shot fine-tuning models. This number represents a good trade-off between the computational cost of the methods and the quality of samples. Running the training for full number of epochs may result in better selection of samples, but we consider such setting to not be representative. In addition, we use the same number of epochs for each strategy to keep the comparison unbiased and representative. In addition to searching for optimal number of epochs, we search for an optimal set of hyperparameters for the methods at the same time (such as the learning rate, the optimiser used or batch size). We use the Adam optimiser with Cosine Annealing learning rate scheduler, with learning rate of 0.01 for image data and 0.0001 for text data and a batch size of 64.
To keep the results as comparable and unbiased as possible, each selection strategy uses the same base model as the one used in the main experiments, i.e., ResNet-18 (He et al., 2016) pretrained on Imagenet dataset (Deng et al., 2009) for image data and pretrained BERT-base (Devlin et al., 2019) for text data in case of gradient few-shot learning, and the same large language models (Mistral (Jiang et al., 2023) and Zephyr (Tunstall et al., 2023)), but with only their last classification layer being learned, for in-context learning. In case the selection strategy requires sample features, we use the output of last fully-connected layer (the input of classification layer) as the feature vector similar to (Guo et al., 2022).
We also evaluate the sample selection based on Similarity and Diversity as it represents popular approach for in-context learning. To select the sample, we use the cosine similarity between the feature representation (obtained from the penultimate layer) of the samples. To select the samples for gradient few-shot learning, we first randomly select one sample for each class. Afterwards, we expand the set of selected samples by iteratively adding the most similar (for Similarity) or most dissimilar (for Diversity) sample to the average feature representation of the set of selected samples for each class (i.e., if we already have 5 selected samples in the specific class, we calculate the similarity to the average vector feature representation of these 5 samples). For in-context learning, we instead the most similar (for Similarity) or the most dissimilar (for Diversity) samples for each test sample as is common for this approach (Sun et al., 2023).
Finally, we also use the in-context learning specific LENS method (Li & Qiu, 2023) as a baseline. We use the same models (Mistral and Zephyr) for the selection of samples and run a hyperparameter search informed by the recommendations of the authors to an optimal setup for the search. In addition, we use the modified version of their source code for running the method.
For the core-set selection strategies, we use the modified implementation of the methods released by (Guo et al., 2022) and its extension provided by (Park et al., 2022) for active learning methods. For the Cartography strategy, we use the modified implementation released by the original authors (Swayamdipta et al., 2020).
Appendix D Experimental Setup: Further Details
The experiments and evaluation of different sample selection strategies are done on the within-domain few-shot learning, where the classes in training, validation and testing datasets all come from the same domain (e.g., from the same dataset). Each evaluation run is done on 600 randomly selected tasks for the gradient few-shot learning approaches and 300 tasks for the in-context learning approaches (due to high computation costs of LLMs). Each tasks is a 5-way 5-shot (or -shot for experiments where we change the number of shots) classification task evaluated on 16 test samples per class, following the standard few-shot learning methodology and the evaluation from (Ullah et al., 2022). The performance of a single evaluation run is reported as the mean accuracy across these tasks and the different splits of data (we use 5 splits of data as mentioned in the main part of the paper). These scores are then used to report the findings in the main experiments, either as mean and standard deviation over multiple repeated runs of the specific sample selection strategy or over the different models and datasets. As such, the standard deviation of different data splits or different tasks is not reported. This is also the reason why we can observe zero standard deviation on some of the sample selection strategies.
For the gradient few-shot learning approaches we use the ResNet-18 (He et al., 2016) for image data, pretrained on Imagenet dataset (Deng et al., 2009), and a pretrained BERT-base (Devlin et al., 2019) model for the text data as base models. All the models are further trained on each dataset using the different few-shot learning approaches, namely the Prototypical Networks (Snell et al., 2017), Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) and Few-Shot Fine-Tuning (Chen et al., 2019). In case of Few-Shot Fine-Tuning, the further training is done on the concatenation of all the meta-training classes and their corresponding samples, while at the meta-test time, only the last layer is replaced and fine-tuned on the specific task with 5 classes. The Prototypical Networks and MAML follow the typical few-shot training setting, training on large number of 5-way few-shot tasks. The meta-learning approaches on image data are trained on 15 000 tasks and the Few-Shot Fine-Tuning approach is trained on 15 000 randomly sampled batches of size 16. On text data, the number of sampled tasks and batches is increased to 25 000. In addition, the MAML approaches uses 5 training iteration and 10 testing iterations on image data and 10 training and 15 testing on text data, while Prototypical Networks use only 1 in both cases. Finally, all approaches use the Adam optimiser with 0.001 (meta) learning rate and 0.01 base learning rate. The full training process is repeated 3 times and the best performing model, selected using a separate validation dataset, is selected for use for evaluation.
| Datasets | Prompt Format | ||||
| News classification |
| Intent classification |
| Dataset | Verbaliser | ||||
| 20 News Group |
{IBM, Middle East Politics, Windows XP, Motorcycles, Medicine, For Sale, Religion, MS Windows, Baseball, Auto, Hockey, Mac, Graphics, Christianity, Guns, Electronics, Space, Crypto, Atheism, Politics} |
| News Category |
{Politics, World News, Parenting, Money, Wellness, Business, Weddings, Entertainment, Impact, Black Voices, Queer Voices, Crime, Divorce, Food and Drink, Worldpost, Parents, Travel, The Worldpost, Healthy Living, Taste, Media, Culture and Arts, Style, Weird News, Style and Beauty, Comedy, Home and Living, Sports, Environment, Education, Good News, Fifty, Women, Science, Arts and Culture, US News, Green, Tech, Religion, Latino Voices, College, Arts} |
| ATIS |
{Flight, Abbreviation, City, Airfare, Ground Service, Ground Fare, Airline, Flight Number, Aircraft, Distance, Capacity, Flight Time, Quantity, Airport} |
{Road Condition, Departure, Duration, Event, Arrival, Directions, Distance, Traffic, Update, Combine, Route, Location} |
| HWU-64 |
{Set Alarm, Definition, Play Music, Set Calendar, Play Radio, Confirm, Quirky, Send Email, Currency, Like, Social Query, News, Neutral, Stop, Calendar Query, Alarm Query, Recommend Movie, Play Audiobook, Weather Query, Praise, Social Post, Play Podcast, Negate, Recipe, Affirm, Remove Calendar, Recommend Location, Remove List, Explain, Email Query, Maths, Stock, Transport Query, Order Takeaway, Datetime Query, Repeat, Factoid, Taxi, Add List, Add Contact, Recommend Event, Mute Volume, List Query, Ticket, Convert Datetime, Joke, Remove Alarm, Traffic, Volume Up, Takeaway Query, Play Game, Contact Query, Music Query, Volume Other, Volume Down, Music Setting, Greet, Dislike} |
| SNIPS |
{Playlist, Weather, Event, Musing, Creative Work, Rate Book, Book Restaurant} |
For the in-context learning approaches, we use instruction tuned Mistral-7B (Jiang et al., 2023) and Zephyr-7B (Tunstall et al., 2023) large language models in the 4-bit quantised mode. This significantly reduces the computation costs required, while having low impact on the performance and almost no impact the impact of sample selection. We perform no training for these models (except in case of sample selection, which is covered in Appendix C). Instead we run a basic prompt engineering to find the best performing template for each model. In the end, we opted for the simple template that performed the best on average for each model and each dataset, which is illustrated in Table 3, along with the verbaliser (i.e., the possible outputs of the model that are mapped to the classes in the dataset). Finally, the in-context learning models are set to provide deterministic outputs (controlling the temperature and disabling sampling) and set to generate maximum of 10 tokens. In case where multiple words that can be mapped to a dataset class are generated, the output is treated as incorrect as the model is just hallucinating and not working as intended (this behaviour was mainly observed when using the models in zero-shot or 50-shot setting, where the model either does not understand the task or the context is not sufficient for it to follow the instructions).
All the hyperparameters for all models were determined using a hyperparameter search on each data modality and model. For the experiments, we use a modified implementation provided by the MetaAlbum (Ullah et al., 2022), extended to further approaches, models and datasets. In addition, we also the hyperparameters used in the paper as a starting point for the hyperparameter search.
Appendix E Comparison of Sample Selection Impact Between 5-Shot and 10-Shot Setting
To better explore the effect of increasing number of selected shots on the sample selection strategies, we repeat the evaluation of all the selection strategies in the 10-shot setting (using the same setup and methodology) and compare the results to 5-shot setting. We provide the comparison of distribution of results, i.e., the difference to Classic selection, in Figure 5.
Overall, we observe an increase in the overall performance of the models for all the selection strategies (with the exception of some bad performing ones) as well as the Classic selection. At the same time, we observe lower difference to the Classic selection for all the evaluated selection strategies. This finding holds for all strategies across all models and few-shot learning approaches.
In addition, we also compare the results using Wilcoxon signed-rank test, between the results of 5-shot and 10-shot setting. The resulting p-value () indicates that the different to Classic selection is more significant in case of 5-shot setting.
All in all, these experiments further reinforce the findings from comparison of ACSESS and random selection, where we observed lower impact of sample selection on larger number of shots.
Appendix F Dataset Specific Results and Visualisations
In this appendix, we present the full results for each combination of dataset and model we use in the experiments. The results are presented using both figures, showing the difference to Classic selection, as well as tables, showing the overall performance of the model on the specific dataset, the standard deviation from multiple runs of the selection strategy and the difference to the Classic selection (while also showing the results for Classic selection). Following is the mapping for different datasets, each presenting results for all the models:
| LR_AM.DOG | ProtoNet | MAML | Fine-Tuning | |||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||
| Random | ||||||
| Similarity | ||||||
| Diversity | ||||||
| Entropy | ||||||
| Margin | ||||||
| Least Confidence | ||||||
| Loss | ||||||
| CAL | ||||||
| CRAIG | ||||||
| DeepFool | ||||||
| Forgetting | ||||||
| Glister | ||||||
| GraNd | ||||||
| Herding | ||||||
| KCenter | ||||||
| Graph-Cut | ||||||
| CartographyEasy | ||||||
| CartographyAmbiguous | ||||||
| CartographyHard | ||||||
| CartographyEasy+Ambig. | ||||||
| ACSESSUniform | ||||||
| ACSESSWeighted | ||||||
| ACSESSWith Random | ||||||
| LR_AM.AWA | ProtoNet | MAML | Fine-Tuning | |||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||
| Random | ||||||
| Similarity | ||||||
| Diversity | ||||||
| Entropy | ||||||
| Margin | ||||||
| Least Confidence | ||||||
| Loss | ||||||
| CAL | ||||||
| CRAIG | ||||||
| DeepFool | ||||||
| Forgetting | ||||||
| Glister | ||||||
| GraNd | ||||||
| Herding | ||||||
| KCenter | ||||||
| Graph-Cut | ||||||
| CartographyEasy | ||||||
| CartographyAmbiguous | ||||||
| CartographyHard | ||||||
| CartographyEasy+Ambig. | ||||||
| ACSESSUniform | ||||||
| ACSESSWeighted | ||||||
| ACSESSWith Random | ||||||
| VCL.APL | ProtoNet | MAML | Fine-Tuning | |||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||
| Random | ||||||
| Similarity | ||||||
| Diversity | ||||||
| Entropy | ||||||
| Margin | ||||||
| Least Confidence | ||||||
| Loss | ||||||
| CAL | ||||||
| CRAIG | ||||||
| DeepFool | ||||||
| Forgetting | ||||||
| Glister | ||||||
| GraNd | ||||||
| Herding | ||||||
| KCenter | ||||||
| Graph-Cut | ||||||
| CartographyEasy | ||||||
| CartographyAmbiguous | ||||||
| CartographyHard | ||||||
| CartographyEasy+Ambig. | ||||||
| ACSESSUniform | ||||||
| ACSESSWeighted | ||||||
| ACSESSWith Random | ||||||
| HUM_ACT.ACT_410 | ProtoNet | MAML | Fine-Tuning | |||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||
| Random | ||||||
| Similarity | ||||||
| Diversity | ||||||
| Entropy | ||||||
| Margin | ||||||
| Least Confidence | ||||||
| Loss | ||||||
| CAL | ||||||
| CRAIG | ||||||
| DeepFool | ||||||
| Forgetting | ||||||
| Glister | ||||||
| GraNd | ||||||
| Herding | ||||||
| KCenter | ||||||
| Graph-Cut | ||||||
| CartographyEasy | ||||||
| CartographyAmbiguous | ||||||
| CartographyHard | ||||||
| CartographyEasy+Ambig. | ||||||
| ACSESSUniform | ||||||
| ACSESSWeighted | ||||||
| ACSESSWith Random | ||||||
| MNF.TEX_DTD | ProtoNet | MAML | Fine-Tuning | |||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||
| Random | ||||||
| Similarity | ||||||
| Diversity | ||||||
| Entropy | ||||||
| Margin | ||||||
| Least Confidence | ||||||
| Loss | ||||||
| CAL | ||||||
| CRAIG | ||||||
| DeepFool | ||||||
| Forgetting | ||||||
| Glister | ||||||
| GraNd | ||||||
| Herding | ||||||
| KCenter | ||||||
| Graph-Cut | ||||||
| CartographyEasy | ||||||
| CartographyAmbiguous | ||||||
| CartographyHard | ||||||
| CartographyEasy+Ambig. | ||||||
| ACSESSUniform | ||||||
| ACSESSWeighted | ||||||
| ACSESSWith Random | ||||||
| MCR.PRT | ProtoNet | MAML | Fine-Tuning | |||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||
| Random | ||||||
| Similarity | ||||||
| Diversity | ||||||
| Entropy | ||||||
| Margin | ||||||
| Least Confidence | ||||||
| Loss | ||||||
| CAL | ||||||
| CRAIG | ||||||
| DeepFool | ||||||
| Forgetting | ||||||
| Glister | ||||||
| GraNd | ||||||
| Herding | ||||||
| KCenter | ||||||
| Graph-Cut | ||||||
| CartographyEasy | ||||||
| CartographyAmbiguous | ||||||
| CartographyHard | ||||||
| CartographyEasy+Ambig. | ||||||
| ACSESSUniform | ||||||
| ACSESSWeighted | ||||||
| ACSESSWith Random | ||||||
| MCR.PNU | ProtoNet | MAML | Fine-Tuning | |||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||
| Random | ||||||
| Similarity | ||||||
| Diversity | ||||||
| Entropy | ||||||
| Margin | ||||||
| Least Confidence | ||||||
| Loss | ||||||
| CAL | ||||||
| CRAIG | ||||||
| DeepFool | ||||||
| Forgetting | ||||||
| Glister | ||||||
| GraNd | ||||||
| Herding | ||||||
| KCenter | ||||||
| Graph-Cut | ||||||
| CartographyEasy | ||||||
| CartographyAmbiguous | ||||||
| CartographyHard | ||||||
| CartographyEasy+Ambig. | ||||||
| ACSESSUniform | ||||||
| ACSESSWeighted | ||||||
| ACSESSWith Random | ||||||
| PLT.FLW | ProtoNet | MAML | Fine-Tuning | |||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||
| Random | ||||||
| Similarity | ||||||
| Diversity | ||||||
| Entropy | ||||||
| Margin | ||||||
| Least Confidence | ||||||
| Loss | ||||||
| CAL | ||||||
| CRAIG | ||||||
| DeepFool | ||||||
| Forgetting | ||||||
| Glister | ||||||
| GraNd | ||||||
| Herding | ||||||
| KCenter | ||||||
| Graph-Cut | ||||||
| CartographyEasy | ||||||
| CartographyAmbiguous | ||||||
| CartographyHard | ||||||
| CartographyEasy+Ambig. | ||||||
| ACSESSUniform | ||||||
| ACSESSWeighted | ||||||
| ACSESSWith Random | ||||||
| 20 News Group | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||||||
| Random | ||||||||||
| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
| Similarity | ||||||||||
| Diversity | ||||||||||
| Entropy | ||||||||||
| Margin | ||||||||||
| Least Confidence | ||||||||||
| Loss | ||||||||||
| CAL | ||||||||||
| CRAIG | ||||||||||
| DeepFool | ||||||||||
| Forgetting | ||||||||||
| Glister | ||||||||||
| GraNd | ||||||||||
| Herding | ||||||||||
| KCenter | ||||||||||
| Graph-Cut | ||||||||||
| CartographyEasy | ||||||||||
| CartographyAmbiguous | ||||||||||
| CartographyHard | ||||||||||
| CartographyEasy+Ambig. | ||||||||||
| ACSESSUniform | ||||||||||
| ACSESSWeighted | ||||||||||
| ACSESSWith Random | ||||||||||
| News Category | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||||||
| Random | ||||||||||
| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
| Similarity | ||||||||||
| Diversity | ||||||||||
| Entropy | ||||||||||
| Margin | ||||||||||
| Least Confidence | ||||||||||
| Loss | ||||||||||
| CAL | ||||||||||
| CRAIG | ||||||||||
| DeepFool | ||||||||||
| Forgetting | ||||||||||
| Glister | ||||||||||
| GraNd | ||||||||||
| Herding | ||||||||||
| KCenter | ||||||||||
| Graph-Cut | ||||||||||
| CartographyEasy | ||||||||||
| CartographyAmbiguous | ||||||||||
| CartographyHard | ||||||||||
| CartographyEasy+Ambig. | ||||||||||
| ACSESSUniform | ||||||||||
| ACSESSWeighted | ||||||||||
| ACSESSWith Random | ||||||||||
| ATIS | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||||||
| Random | ||||||||||
| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
| Similarity | ||||||||||
| Diversity | ||||||||||
| Entropy | ||||||||||
| Margin | ||||||||||
| Least Confidence | ||||||||||
| Loss | ||||||||||
| CAL | ||||||||||
| CRAIG | ||||||||||
| DeepFool | ||||||||||
| Forgetting | ||||||||||
| Glister | ||||||||||
| GraNd | ||||||||||
| Herding | ||||||||||
| KCenter | ||||||||||
| Graph-Cut | ||||||||||
| CartographyEasy | ||||||||||
| CartographyAmbiguous | ||||||||||
| CartographyHard | ||||||||||
| CartographyEasy+Ambig. | ||||||||||
| ACSESSUniform | ||||||||||
| ACSESSWeighted | ||||||||||
| ACSESSWith Random | ||||||||||
| ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | ||||||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||||||
| Random | ||||||||||
| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
| Similarity | ||||||||||
| Diversity | ||||||||||
| Entropy | ||||||||||
| Margin | ||||||||||
| Least Confidence | ||||||||||
| Loss | ||||||||||
| CAL | ||||||||||
| CRAIG | ||||||||||
| DeepFool | ||||||||||
| Forgetting | ||||||||||
| Glister | ||||||||||
| GraNd | ||||||||||
| Herding | ||||||||||
| KCenter | ||||||||||
| Graph-Cut | ||||||||||
| CartographyEasy | ||||||||||
| CartographyAmbiguous | ||||||||||
| CartographyHard | ||||||||||
| CartographyEasy+Ambig. | ||||||||||
| ACSESSUniform | ||||||||||
| ACSESSWeighted | ||||||||||
| ACSESSWith Random | ||||||||||
| HWU-64 | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||||||
| Random | ||||||||||
| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
| Similarity | ||||||||||
| Diversity | ||||||||||
| Entropy | ||||||||||
| Margin | ||||||||||
| Least Confidence | ||||||||||
| Loss | ||||||||||
| CAL | ||||||||||
| CRAIG | ||||||||||
| DeepFool | ||||||||||
| Forgetting | ||||||||||
| Glister | ||||||||||
| GraNd | ||||||||||
| Herding | ||||||||||
| KCenter | ||||||||||
| Graph-Cut | ||||||||||
| CartographyEasy | ||||||||||
| CartographyAmbiguous | ||||||||||
| CartographyHard | ||||||||||
| CartographyEasy+Ambig. | ||||||||||
| ACSESSUniform | ||||||||||
| ACSESSWeighted | ||||||||||
| ACSESSWith Random | ||||||||||
| SNIPS | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
| Classic | ||||||||||
| Random | ||||||||||
| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
| Similarity | ||||||||||
| Diversity | ||||||||||
| Entropy | ||||||||||
| Margin | ||||||||||
| Least Confidence | ||||||||||
| Loss | ||||||||||
| CAL | ||||||||||
| CRAIG | ||||||||||
| DeepFool | ||||||||||
| Forgetting | ||||||||||
| Glister | ||||||||||
| GraNd | ||||||||||
| Herding | ||||||||||
| KCenter | ||||||||||
| Graph-Cut | ||||||||||
| CartographyEasy | ||||||||||
| CartographyAmbiguous | ||||||||||
| CartographyHard | ||||||||||
| CartographyEasy+Ambig. | ||||||||||
| ACSESSUniform | ||||||||||
| ACSESSWeighted | ||||||||||
| ACSESSWith Random | ||||||||||

### 17 Apr 2026 v2

@@
 Abstract
-In few-shot learning, such as meta-learning, few-shot fine-tuning or in-context learning, the limited number of samples used to train a model have a significant impact on the overall success.
-Although a large number of sample selection strategies exist, their impact on the performance of few-shot learning is not extensively known, as most of them have been so far evaluated in typical supervised settings only.
-In this paper, we thoroughly investigate the impact of 20 sample selection strategies on the performance of 5 few-shot learning approaches over 8 image and 6 text datasets.
-In addition, we propose a new method for automatic combination of sample selection strategies (ACSESS) that leverages the strengths and complementary information of the individual strategies.
-The experimental results show that our method consistently outperforms the individual selection strategies, as well as the recently proposed method for selecting support examples for in-context learning.
-We also show a strong modality, dataset and approach dependence for the majority of strategies as well as their dependence on the number of shots – demonstrating that the sample selection strategies play a significant role for lower number of shots, but regresses to random selection at higher number of shots.
-[…]
+In few-shot learning, the selection of samples has a significant impact on the performance of the model.
+While effective sample selection strategies are well-established in supervised settings, research on large language models largely overlooks them, favouring strategies specifically tailored to individual in-context learning settings.
+In this paper, we propose a new method for Automatic Combination of SamplE Selection Strategies (ACSESS) to leverage the strengths and complementarity of various well-established selection objectives.
+We investigate and compare the impact of 23 sample selection strategies on the performance of 5 in-context learning models and 3 few-shot learning approaches (meta-learning, few-shot fine-tuning) over 6 text and 8 image datasets.
+The experimental results show that the combination of strategies through the ACSESS method consistently outperforms all individual selection strategies and performs on par or exceeds the in-context learning specific baselines.
+Lastly, we demonstrate that sample selection remains effective even on smaller datasets, yielding the greatest benefits when only a few shots are selected, while its advantage diminishes as the number of shots increases.
+Automatic Combination of Sample Selection Strategies for Few-Shot Learning
+Branislav Pecher♠†, Ivan Srba†, Maria Bielikova†, Joaquin Vanschoren‡ ♠ Faculty of Information Technology, Brno University of Technology, Brno, Czechia † Kempelen Institute of Intelligent Technologies, Bratislava, Slovakia ‡ Eindhoven University of Technology, Eindhoven, Netherlands {branislav.pecher, ivan.srba, maria.bielikova}@kinit.sk, j.vanschoren@tue.nl
 1 Introduction
-Many domains are characterised by a labelled data scarcity due to data collection and annotation costs and privacy, making the use of typical deep learning unfeasible.
-To address these problems, few-shot learning aims to train or adapt machine learning models to new tasks using only few labelled samples per class (Song et al., 2023).
-There are two main categories of few-shot learning (more details regarding their difference are included in Appendix B): 1) gradient-based, which includes meta-learning and fine-tuning, where the knowledge from relevant datasets is transferred to new problem using gradient-based training, similar to supervised learning (Song et al., 2023; Chen et al., 2019; Vanschoren, 2018; Hospedales et al., 2021); and 2) few-shot learning via in-context learning, a new paradigm, where a pretrained large language model is conditioned on few examples to generate the prediction without any parameter updates (Dong et al., 2022; Liu et al., 2023).
-As the number of utilised labelled samples is small, the role of individual samples is all the more important.
-Based on the choice of samples for in-context learning, the observed performance can vary between state-of-the-art and random predictions, indicating a strong sensitivity to the sample selection (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022).
-Even though a lot of research is dedicated to sample selection, such as active learning (Ren et al., 2021) or core-set selection (Guo et al., 2022), the most common approach in few-shot learning is a simple random selection.
-The exception is in-context learning, where a large number of sample selection strategies have been proposed recently, selecting samples according to their properties such as similarity, diversity, informativeness or quality (Li & Qiu, 2023; Zhang et al., 2022; Chang & Jia, 2023).
-However, the impact of sample selection strategies on in-context learning or gradient-based few-shot learning remains largely under-explored.
-Therefore, the goal in this paper is to investigate the impact of the sample selection strategies on few-shot image and text classification tasks across a wide range of datasets.
-In addition, as the different strategies mostly focus on selecting samples based on a single property (e.g., how often the samples are forgotten (Toneva et al., 2018), how easy they are to learn (Swayamdipta et al., 2020; Zhang & Plank, 2021) or how informative they are (Ye et al., 2023b; Liu & Wang, 2023)), we ask whether the selection based on their combination would not lead to better results.
-To answer this question and to leverage the strengths of different sample selection strategies, we propose ACSESS, an effective method for Automatic Combination of SamplE Selection Strategies.
-The proposed method first identifies a subset of relevant strategies that can improve the overall success of few-shot learning.
-Afterwards, the identified strategies are combined together using weighting based on their expected contribution to find samples that can provide the most benefit.
-We summarise our contributions as follows111To support replicability and extension of our results, we openly publish the source code of our experiments at https://anonymous.4open.science/r/ACSESS:
-We evaluate the impact of 20 sample selection strategies on 5 few-shot learning approaches, including meta-learning, few-shot fine-tuning and in-context learning, across 8 image and 6 text datasets.
-Experimental results show the selection strategies often lead to non-negligible improvements in performance (up to percentage points for gradient few-shot learning or percentage points for in-context learning), while showing a strong approach and dataset dependence.
-We propose a method for automatic combination of sample selection strategies to leverage their strengths and identify samples with complementary properties required for the success of few-shot learning.
-Experimental results show the method consistently leads to performance improvement (up to percentage points for gradient few-shot learning or percentage points for in-context learning), outperforming all investigated strategies, including a recent in-context specific strategy LENS (Li & Qiu, 2023).
-We further analyse how the impact of sample selection changes when increasing the number of shots.
-We find following key insights: 1) sample selection has significantly higher impact on lower number of shots; 2) at higher number of shots (30-40 on average), the impact of sample selection is negligible as all strategies regress to random selection; and 3) after a certain point, the boost in performance from increasing the number of shots is negligible (50-shot for gradient few-shot learning; 20-shot for in-context learning with strong dependence on input text size and model context size).
+Many domains are characterised by a labelled data scarcity due to data collection/annotation costs or privacy considerations, making the training of typical deep learning models unfeasible.
+Few-shot learning addresses the challenge of adapting models to new tasks when only a handful of labelled samples are available (Song et al., 2023).
+Two main approaches exist: 1) in-context learning, where a pretrained large language model is conditioned on few examples without any parameter updates (Dong et al., 2022; Liu et al., 2023); and 2) gradient-based few-shot learning, which includes meta-learning and fine-tuning, where the knowledge from relevant datasets is transferred to a new problem using gradient-based training (Song et al., 2023; Chen et al., 2019; Vanschoren, 2018; Hospedales et al., 2021).
+More details regarding their differences are included in Appendix J.
+The choice of which sample to use is critical in both approaches, as performance can vary drastically depending on sample quality.
+Existing studies in different settings, such as in LLM alignment (Zhou et al., 2024), have shown that curating a smaller subset of high-quality samples can lead to better performance than training on the full set of available samples.
+This holds for few-shot learning as well, especially for noisy or unbalanced datasets or for in-context learning, which is known to be sensitive to sample choice (Pecher et al., 2024; Agarwal et al., 2021; Köksal et al., 2022).
+Many works have already proposed diverse sample selection strategies to identify an optimal set of samples.
+Such a sample selection strategy follows its specific objective by considering one or few properties of available samples, e.g., to maximise similarity, diversity, informativeness or quality of selected samples (Li and Qiu, 2023; Zhang et al., 2022; Chang and Jia, 2023).
+With the popularisation of large language models (LLMs), many new selection strategies are proposed for in-context learning.
+However, they are often tailored only for specific settings, limiting their applicability.
+At the same time, the well-established and long-standing selection strategies that were shown to perform well in supervised settings, such as active learning, are mostly ignored as part of LLM sample selection (Pecher et al., 2024; Albalak et al., 2024).
+In this work, we demonstrate that well-established strategies can match or exceed the performance of in-context learning specific strategies.
+To achieve this, we automatically and optimally combine the sample selection objectives used by well-established sample selection strategies to identify samples with complementary properties.
+In order to leverage the strengths of different sample selection strategies, we propose ACSESS, an effective method for Automatic Combination of SamplE Selection Strategies.
+First, a subset of relevant strategies that can improve the overall success of few-shot learning is identified.
+Afterwards, the selection objectives from the identified strategies are combined together, using weighting based on their expected contribution, in order to identify the most informative and high-quality samples that can provide the most benefit.
+Our contributions and findings are as follows111To support replicability and extension of our results, we openly publish the source code of our experiments at https://github.com/kinit-sk/ACSESS:
+We provide the first large-scale study of 23 sample selection strategies across 5 in-context learning models and 3 few-shot learning approaches, including meta-learning and few-shot fine-tuning, across 6 text and 8 image datasets.
+We show that sample selection can yield consistent gains (up to and percentage points for in-context learning and gradient few-shot learning, respectively), though the effects depend heavily on dataset and approach.
+We propose ACSESS, a method that automatically combines and weights sample selection objectives to leverage their strengths and identify samples with complementary properties (i.e., informativeness, representativeness and learnability).
+Our results show that the combination of well-established strategies identified by the ACSESS method consistently leads to performance on par or better than the existing in-context learning specific baselines.
+Through ablation studies, we find following key insights: 1) sample selection has higher impact when the number of shots is low or when using noisy datasets; 2) at higher number of shots (30-40 on average), the impact of sample selection is negligible as all strategies regress to random selection; 3) after a certain point, the boost in performance from using more shots becomes negligible (30 shots for in-context learning; 50-shot for gradient few-shot learning); 4) sample selection is beneficial even for small dataset sizes (achieving similar performance when selecting from only 25% for in-context learning or 10% of the dataset for gradient few-shot learning).
 2 Related Work
-A lot of focus is dedicated to sample selection: 1) to reduce the computation costs without the loss of performance, by distilling dataset to lower number of samples (Yu et al., 2023); 2) selecting a core set of a dataset with similar properties and same performance as the full set (Guo et al., 2022); or 3) to reduce annotation costs using active learning (Ren et al., 2021).
-A specific research focus is dedicated to selecting good performing samples for in-context learning, as the overall performance was found to be sensitive to this choice, leading to large variability in results (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022).
-However, the adoption of sample selection in few-shot learning is limited to active learning strategies and in-context learning.
-The sample selection strategies can be divided into two groups.
-The first, more prevalent, uses heuristics and unsupervised metrics to rate and select samples according to specific property that serves as estimate of their potential to increase the model’s performance.
-The most popular approach is selecting samples based on the similarity to the test sample for in-context learning.
-The methods differ in employed: 1) similarity measure, such as cosine/Euclidean distance (Liu et al., 2022; An et al., 2023; Zemlyanskiy et al., 2022; Gupta et al., 2022; Pasupat et al., 2021), BM-25 (Wang et al., 2022; Nashid et al., 2023) or their combination (Agrawal et al., 2023; Gao et al., 2021); and 2) representation, such as simple TF-IDF (Agrawal et al., 2023; Wang et al., 2022; Zemlyanskiy et al., 2022), word embeddings from basic or further fine-tuned transformers (Liu et al., 2022; Gao et al., 2021; Gupta et al., 2022; Pasupat et al., 2021), or representation generated by prompting a large language model (An et al., 2023; Nashid et al., 2023).
-Besides similarity, the samples are often selected based on their informativeness or the uncertainty.
-The active learning strategies are most common (Köksal et al., 2022; Schröder et al., 2022; Margatina et al., 2021; Park et al., 2022; Mavromatis et al., 2023), but also other approaches are utilised, such as using the bias of the samples (Ma et al., 2023).
-Other methods define a notion of quality for each sample, either by prompting a large language model to rate the samples (Shin et al., 2021) or observing how the inclusion or removal of the sample affects the performance (Maronikolakis et al., 2023; Nguyen & Wong, 2023).
-When considering representativeness, the samples are selected using core-set methods that are designed to find a subset of samples with properties that are representative of the full training set (Guo et al., 2022; Killamsetty et al., 2021; Mirzasoleiman et al., 2020; Paul et al., 2021; Coleman et al., 2020).
-Learnability metric either determines how easy it is to learn the specific sample (Swayamdipta et al., 2020; Zhang & Plank, 2021), or how often the sample is forgotten after being learned (Toneva et al., 2018).
-Finally, some approaches balance multiple properties at the same time.
-Often, the similarity of samples is balanced with their diversity (Wu et al., 2023; Qin et al., 2023) or the informativeness of the samples is balanced with their representativeness (Su et al., 2022; Levy et al., 2023; Ye et al., 2023b; Liu & Wang, 2023).
-In specific cases, the search is done in two step fashion, such as finding set of informative samples and then using diversity guided search to improve this set (Li & Qiu, 2023).
-The second set of approaches leverages training to select the samples.
-One possibility is to train a sample retriever, such as a separate scoring model (Rubin et al., 2022; Luo et al., 2023; Li et al., 2023; Wang et al., 2023; Aimen et al., 2023) or using determinantal point process (Yang et al., 2023; Ye et al., 2023a).
-Another set of approaches leverages reinforcement learning to select good samples (Zhang et al., 2022; Shum et al., 2023; Scarlatos & Lan, 2023).
-Finally, the Datamodels approach trains a linear model to predict the performance gain of a set of samples to select the subset that would lead to the highest possible performance increase (Ilyas et al., 2022; Chang & Jia, 2023; Jundi & Lapesa, 2022; Vilar et al., 2023).
-In this work, we investigate the impact of the existing strategies on few-shot setting.
-In contrast to (Li & Qiu, 2023), which represents a work closest to ours, we explore multiple few-shot learning approaches not only in-context learning.
-Our proposed method is inspired by the Datamodels approach (Ilyas et al., 2022), but performs the selection at the level of strategies instead of samples.
-3 Sample Selection Strategies
-3.1 Single-Property Strategies
-| Strategy | Description | |
-Informativeness |
-Similarity |
-Most similar samples (popular for in-context learning). |
-Diversity |
-Most diverse/dissimilar samples. |
-Entropy (Park et al., 2022) |
-Active Learning; Highest entropy over class probabilities. |
-Margin (Park et al., 2022) |
-Active Learning; Lowest difference in probability between top 2 classes. |
-Least Confidence (Park et al., 2022) |
-Active Learning; Top class with lowest probability assigned. |
-Loss (Park et al., 2022) |
-Active Learning; Highest loss. |
-Contrastive Active Learning (CAL) (Margatina et al., 2021) |
-Core-Set; Predictive likelihood that diverges most from the neighbourhood as determined by KL divergence. |
-GraNd (Paul et al., 2021) |
-Core-Set; Highest contribution to decline of training loss. |
-Core-Set; Submodularity function that measures diversity and informativeness. |
-Representativeness |
-Core-Set; Minimising distance of subset centre to full dataset. |
-Core-Set; Minimising distance of every sample in subset to full dataset. |
-CRAIG (Mirzasoleiman et al., 2020) |
-Core-Set; Gradients representative of full dataset. |
-Glister (Killamsetty et al., 2021) |
-Core-Set; Bi-level optimisation. |
-Learnability |
-Forgetting (Toneva et al., 2018) |
-Core-Set; How often the samples are forgotten. |
-How easy it is to learn the samples. Due to no obvious consensus, all of easy, ambiguous, hard and combination of easy and ambiguous samples are considered. |
-In this work, we thoroughly investigate the impact of strategies that consider 3 groups of properties: 1) informativeness, or how informative the samples are for the model; 2) representativeness, or how well the subset of samples represents the full dataset; and 3) learnability, or how easy it is to obtain or retain the information contained in the samples.
-We specifically focus on strategies that perform the selection by assigning a score to each sample that is used to sort the samples and choose the best.
-This includes active learning strategies that measure the uncertainty of the model, core-set selection strategies that find subset representative of the full-set and other strategies that use training dynamics.
-The high-level overview of these strategies is presented in Table 1 and their further details are included in Appendix C.
-3.2 ACSESS: Automatic Combination of Strategies
-Even though the different selection strategies can identify samples with specific single property, the use of these identified samples may not have a significant impact on the performance when working in a few-shot setting.
-For example, an informative sample that is a hard to learn or is often forgotten, may not contribute as much, leading to low impact on the performance.
-Instead, we propose to select examples that are characterised by a balance of complementary properties, even when they do not represent the best sample of any given property (e.g., a sample that may be less informative, but is more easily learned or retained).
-To accomplish this, we propose ACSESS, an effective method for Automatic Combination of SamplE Selection Strategies, composed of two stages: 1) identification of a subset of well-performing and relevant strategies that can improve the overall success of few-shot learning; and then 2) combination of the identified strategies using a weighted scheme based on their expected contribution.
-The final set of samples is selected by sorting the samples according to their final score and selecting the ones with the highest score.
+A large body of work is dedicated to selecting high-quality samples for in-context learning, where the overall performance was found to be sensitive to this choice, leading to large variability in results (Pecher et al., 2024; Köksal et al., 2022; Zhang et al., 2022).
+In the supervised setting, the selection is mostly focused on distilling datasets to a lower number of samples (Yu et al., 2023), selecting a core set representative of the full set (Guo et al., 2022), or reducing annotation costs using active learning (Ren et al., 2021).
+However, the impact of these strategies remains underexplored for other few-shot learning approaches.
+Within the majority of the sample selection strategies, the sample selection objective takes into consideration a single property of available samples.
+To this end, various heuristics and unsupervised metrics serve as an estimate of their potential to increase the model’s performance.
+For in-context learning, the most popular approach is selecting samples based on the similarity to the test sample, differing only in employed similarity measure and representation (Liu et al., 2022; An et al., 2023; Zemlyanskiy et al., 2022; Gupta et al., 2022; Pasupat et al., 2021; Wang et al., 2022; Nashid et al., 2023; Agrawal et al., 2023; Gao et al., 2021).
+Besides similarity, the samples are often selected based on their informativeness or the uncertainty, often using active learning strategies (Köksal et al., 2022; Schröder et al., 2022; Margatina et al., 2021; Park et al., 2022; Mavromatis et al., 2023), bias of the samples (Ma et al., 2023) or adversarial training (Agarwal et al., 2021).
+Other methods define a notion of quality for each sample, either by prompting a LLM to rate the samples (Shin et al., 2021) or observing how the inclusion or removal of the sample affects the performance (Maronikolakis et al., 2023; Nguyen and Wong, 2023).
+When considering representativeness, the samples are selected based on how well they are representative of the full dataset (Guo et al., 2022; Killamsetty et al., 2021; Mirzasoleiman et al., 2020; Paul et al., 2021; Coleman et al., 2020).
+Learnability metric either determines how easy it is to learn the specific sample (Swayamdipta et al., 2020; Zhang and Plank, 2021), or how often the sample is forgotten after being learned (Toneva et al., 2018).
+Finally, some approaches balance multiple sample properties at the same time.
+Often, the similarity of samples is combined with their diversity (Wu et al., 2023; Qin et al., 2023) or the informativeness of the samples is balanced with their representativeness (Su et al., 2022; Levy et al., 2023; Ye et al., 2023; Liu and Wang, 2023).
+In specific cases, a two step sample search is done, such as finding set of informative samples and then using diversity guided search to improve this set (Li and Qiu, 2023), or using different uncertainty measures to select the top samples (Diao et al., 2024).
+Newer strategies leverage optimisation to select a single representative set of samples.
+One possibility is to train a sample retriever, such as a separate scoring model (Rubin et al., 2022; Luo et al., 2023; Li et al., 2023; Wang et al., 2023; Aimen et al., 2023) or define a surrogate scorer for modelling the sample’s goodness and use multi-armed bandit to select the best exemplars (Purohit et al., 2024, 2025).
+Another option is to leverage reinforcement learning (Zhang et al., 2022; Shum et al., 2023; Scarlatos and Lan, 2023).
+Finally, the Datamodels approach trains a linear model to predict the performance gain of a set of samples to select the subset that would lead to the highest possible performance increase (Ilyas et al., 2022; Chang and Jia, 2023; Jundi and Lapesa, 2022; Vilar et al., 2023).
+In this work, we investigate whether the combination of objectives from well-established single-property strategies can rival the selection from the strategies tailored specifically to in-context learning.
+Similar to the close works of Li and Qiu (2023); Purohit et al.
+(2024, 2025), we select a single set of samples, but do not focus solely on in-context learning.
+In essence, we complement work by Agarwal et al.
+(2021) by exploring and effectively combining sample properties important for few-shot learning performance.
+Finally, our proposed method is inspired by the Datamodels approach (Ilyas et al., 2022), but performs the selection at the level of strategies instead of samples.
+3 ACSESS: Automatic Combination of Selection Strategies
+Despite the large number of existing sample selection strategies, most remain constrained to a single objective, typically considering only one sample property.
+This narrow focus often leads to suboptimal selection, whereas greater performance gains can be achieved by appropriately combining multiple objectives that capture complementary sample properties.
+For example, the most informative sample that is hard to learn or is often forgotten may not contribute as much for few-shot methods that utilise gradient-based training.
+On the other hand, for in-context learning, the samples on the decision boundary, which are often hard to learn, may provide more benefit.
+Instead of focusing on a single property, we propose to select examples that are characterised by a balance of complementary properties, even when they may not represent the best sample of any given individual property.
+To accomplish this, we propose ACSESS, a method for Automatic Combination of SamplE Selection Strategies taking their complementarity into account.
+The method is composed of three stages: 1) definition of single-property strategies that will be considered for the combination; 2) identification of a subset of well-performing and relevant strategies that can improve the performance of few-shot learning; and then 3) combining the identified strategies to identify the final set of the most beneficial samples.
+To combine selection objectives, the ACSESS method assigns a selection score for each sample, representing its expected contribution towards a higher performance.
+It is calculated as a weighted sum of scores provided by individual objectives, formally , where represents the set of considered strategies, is the weight assigned to the strategy , is the input sample to be scored, and is the score assigned to the sample by strategy normalised to a interval.
+After assigning a score to each sample , samples with the highest score are selected.
+The steps of the method are summarised in Algorithm 1, with further details, including discussion of the computation complexity and its optimisation, in Appendix A.
+Single-Property Strategies
+We focus on well-established strategies that select based on following properties: 1) informativeness, or how informative the samples are for the model; 2) representativeness, or how well the subset of samples represents the full dataset; and 3) learnability, or how easy it is to obtain or retain the information contained in the samples.
+In this work, we specifically focus on strategies that work by assigning a score to each sample, sorting based on this score and selecting the samples with the best score.
+For informativeness, this includes selecting samples based in their Similarity and Diversity, active learning strategies, such as Entropy, Margin, Least Confidence and Loss (Park et al., 2022), and core-set selection including Contrastive Active Learning (CAL) (Margatina et al., 2021), DeepFool (Ducoffe and Precioso, 2018), GraNd (Paul et al., 2021) and Graph-Cut (Iyer and Bilmes, 2013; Iyer et al., 2021).
+For representativeness, we consider Herding (Welling, 2009; Chen et al., 2010), KCenter (Sener and Savarese, 2018; Agarwal et al., 2020), CRAIG (Mirzasoleiman et al., 2020) and Glister (Killamsetty et al., 2021).
+Finally, for learnability, we consider Forgetting (Toneva et al., 2018) and Cartography (Swayamdipta et al., 2020; Zhang and Plank, 2021), selecting all of each, ambiguous, hard and a combination of easy and ambiguous samples (due to no obvious consensus on which are the best).
+More details regarding the strategies and their objectives are included in Appendix B.
 Identifying the Subset of Relevant Strategies
-In the first phase, the aim is to identify relevant strategies that, when combined, can select high quality samples based on their different properties.
-Even though the most optimal combination would be achieved by searching through all the combinations of strategies, such search is infeasible as it requires extensive computation costs.
-Instead, we use three separate methods for selecting strategies, inspired by feature selection in traditional machine learning and the Datamodels approach (Ilyas et al., 2022), specifically: 1) forward strategy selection; 2) backward strategy selection; and 3) datamodels strategy selection.
-The final subset of selection strategies is composed of the ones that are identified by each of the methods as most impactful, to minimise the effect of potential drawbacks of the individual method.
-Forward Selection iteratively adds to the subset the selection strategy that results in the highest performance increase (similarly to forward feature selection).
+In the second phase, the goal is to identify relevant strategies that, when combined, can select high-quality samples based on their complementary properties.
+Even though the optimal combination could be achieved by searching through all the combinations of strategies, such a search is infeasible as it requires extensive computation.
+Instead, we draw inspiration from methods commonly used for feature selection in traditional machine learning – namely the forward and backward selection – and for the first time apply them to the problem of sample selection.
+In addition, we adapt Datamodels, introduced by Ilyas et al.
+(2022) for individual sample selection, to work on the level of strategies instead.
+Such an approach allows us to explore a more diverse set of possible strategy combinations as compared to forward/backward selection.
+All three of the methods are run independently of each other, each producing a separate set of strategies, i.e., , and respectively for forward, backward and datamodels selection.
+The final set of strategies used by our method is constructed as an intersection between these sets, i.e., .
+This allows the method to identify the most impactful strategies (i.e., strategies identified by all of the methods), minimise the potential drawbacks of the individual methods (i.e., exploring also diverse combinations), and guarantee efficiency of the method, by identifying the lowest number of strategies possible.
+Forward Selection iteratively adds strategies that result in the highest performance increase (similarly to forward feature selection).
 Starting with an empty subset, the method alternates between two steps.
-First, for each strategy not included in the subset we determine the increase in performance the strategy would yield if added to the subset (evaluating the performance of samples selected by the combination of already selected subset and the potential new strategy and comparing to old subset).
-Afterwards, the strategy that provides the highest positive increase in performance is added to the subset and the first step is repeated.
-If no added strategy leads to increase of the overall performance, the process ends, resulting in the subset of relevant selection strategies.
-Backward Selection iteratively removes strategies from the subset until the performance no longer increases (similarly to backward feature selection).
-The method works the same way as the forward selection, but starting with all selection strategies and removing them from the subset based on the performance increase such removal leads to.
-Datamodels Selection draws inspiration from the Datamodels (Ilyas et al., 2022; Chang & Jia, 2023), but works on the level of the strategies instead of samples.
-First, a set of 150 random combinations of strategies is created and evaluated.
+First, for each strategy not included in the subset of selected strategies, we determine the increase in performance the strategy would yield if added to the subset.
+We determine the increase by training the model with samples with the highest selection score determined by so far selected strategies (with weights uniformly distributed for all strategies, ).
+Second, the strategy that provides the highest positive increase in performance is added to the subset and the first step is repeated.
+If the addition of any strategy does not lead to a performance increase, the process ends, resulting in the subset of relevant selection strategies () (see Algorithm 2 in Appendix A for more details).
+In essence, the forward selection iteratively adds the best-performing strategies to the set while the performance still increases.
+Backward Selection iteratively removes strategies until the performance no longer increases (similarly to backward feature selection).
+The method works the same way as the forward selection, but starts with all selection strategies and removes them from the subset () based on the performance increase such removal leads to (see Algorithm 3 in Appendix A for more details).
+In essence, the backward selection iteratively removes strategies that lead to lower performance until the performance no longer increases.
+Datamodels Selection draws inspiration from Datamodels (Ilyas et al., 2022; Chang and Jia, 2023).
+First, a set of (in our work ) random combinations of strategies is created and evaluated (using uniform combination).
 The random search is constrained to include each strategy at least 5 times to ensure sufficient coverage of each strategy.
-Afterwards, these combinations are used to create a simple regression dataset, with the difference in performance to the baseline selection (the classic few-shot learning selection) as label, and the multi-label presence vector of strategies as features.
-The dataset is used to train a regularised linear regression, LASSO (Tibshirani, 1996), which zeros out the weights for as many strategies as possible.
-The selection strategies with positive weights in the model represent the subset of relevant strategies.
+Afterwards, these combinations and their performance are used to create a simple regression dataset.
+The difference in performance to the baseline selection (the classic few-shot learning selection) is used as the target.
+As features, we use a multi-label presence vector of strategies, i.e., vector containing only zeros and ones, where 1 is assigned to elements corresponding to strategy that was used in the combination, and 0 for all other ones.
+The dataset is then used to train a regularised linear regression, LASSO (Tibshirani, 1996), which zeros out the weights for as many strategies as possible.
+The selection strategies with positive weights in the model represent the subset of relevant strategies () (see Algorithm 4 in Appendix A for more details).
 Combining the Identified Strategies
-After the relevant selection strategies are identified, they are combined using a weighting scheme.
-Each sample is assigned an importance score by each of the identified strategies.
-As the different strategies may assign scores of different magnitudes, all scores are normalised to zero-one interval.
-The final sample score is obtained as a weighted average of its assigned scores and the samples with highest score is selected.
+After identifying the relevant selection strategies (), a weighted combination of objectives is used to perform the selection. The final set consists of samples with the highest selection score.
 In this paper, we evaluate three separate weighting schemes.
-First, uniform combination, assigns the same weight to each of the identified strategies.
-This represents the simplified version of the ACSESS method, potentially resulting in lower performance impact, but also with lower computation costs as no further optimisation is required.
-This weighting is also used for forward and backward selection.
-Second, weighted combination, assigns a different weight to each strategy, prioritising specific sample properties.
-The final trained weights of the Datamodels linear model are used as the dataset and model specific weights for the different strategies.
-As the weighting scheme is dataset and model specific, it should lead to the highest performance impact, at the cost of requiring more extensive computation.
-Finally, the weighted combination with random selection introduces a random element to the weighted combination, as random selection was identified as a strong baseline that can improve active learning performance (Park et al., 2022).
-Each sample is assigned additional uniformly randomised score which is included in the combination.
-The weight of this score is determined by the Datamodels by including random selection as additional strategy, due to its impact being dataset and model dependent.
+First, uniform combination, assigns the same weight to each of the identified strategies ().
+This represents the simplified version of the ACSESS method that potentially leads to lower performance impact, but also lower computation costs and better transferability, as no further optimisation is required.
+This weighting scheme is also used for the previous step for the forward, backward and datamodels selection.
+Second, weighted combination, assigns a different weight to each strategy objective, prioritising specific sample properties beneficial for the specific few-shot approach.
+As the strategy weights for this weighted combination, we use the final trained weights of the Datamodels linear model.
+As this weighting scheme is dataset and model-specific, it often leads to the highest performance impact, at the cost of requiring more extensive computation.
+For this reason, in this paper, we evaluate the strategy on a more aggregated level in terms of datasets – see Appendix A for details.
+Finally, the weighted combination with random selection extends the weighted combination by introducing a random element to it, as the random selection was identified as a strong baseline (Park et al., 2022; Cegin et al., 2025).
+For this weighting scheme, each sample is assigned an additional uniformly randomised score.
+To determine the weight of this score, the random selection is included as an additional strategy to the Datamodels method.
 4 Experiments
 Datasets.
-The experiments are conducted on image and text classification datasets composed of different tasks with different number of classes.
-For the image datasets, we use 8 randomly chosen datasets from the Meta-Album (Ullah et al., 2022), specifically LR_AM.DOG and LR_AM.AWA for animal classification, VCL.APL for plane classification, HUM_ACT.ACT_410 for classification of human actions, MNF.TEX_DTD for classification of texture, MCR.PRT and MCR.PNU for classification of microscope images, and PLT.FLW for flower classification.
-In addition, we use 6 widely-adopted text datasets: 20 News Group (Lang, 1995) and News Category (Misra, 2022) for news category classification and ATIS (Hemphill et al., 1990), Facebook (Schuster et al., 2019), HWU-64 (Liu et al., 2021) and SNIPS (Coucke et al., 2018) for intent classification.
+We focus on representative few-shot learning datasets composed of different tasks with different numbers of classes and with different characteristics, such as class imbalance or amount of noisy samples or labels.
+Specifically, we use 6 widely-adopted text datasets, composed of tasks with different number of classes and characteristics, specifically: 20 News Group (Lang, 1995) and News Category (Misra, 2022) for news category classification and ATIS (Hemphill et al., 1990), Facebook (Schuster et al., 2019), HWU-64 (Liu et al., 2021) and SNIPS (Coucke et al., 2018) for intent classification.
+For each dataset, we select a subset of up to 200 labelled samples per class to simulate the limited data setting if needed (the impact of smaller dataset sizes is analysed in Appendix I).
+Furthermore, to evaluate the generalisability, we repeat the experiments on image data as well (see Appendix E for details).
 Few-shot Learning Approaches.
-The evaluation is done in a 5-way 5-shot setting using the gradient few-shot learning on both the image and text datasets and the in-context learning on the text datasets.
-For gradient few-shot learning, we use meta-learning approaches, specifically Prototypical Networks (Snell et al., 2017) and MAML (Finn et al., 2017), and Few-Shot Fine-Tuning (Chen et al., 2019).
-For in-context learning, we use the Mistral-7B (Jiang et al., 2023) and Zephyr-7B (Tunstall et al., 2023) models.
+The evaluation is done in a 5-way 5-shot setting using in-context learning on the text datasets and the gradient few-shot learning on both the text and image datasets.
+For in-context learning, we use the Mistral-7B (Jiang et al., 2023), Zephyr-7B (Tunstall et al., 2023), LLaMA-3.1-8B (Grattafiori et al., 2024), Gemma-3-4B (Team, 2025) and Qwen-2.5-7B (Team, 2024) instruction-tuned large language models.
+For gradient few-shot learning, we use representative and widely-used few-shot learning approaches, specifically 1) meta-learning approaches Prototypical Networks (Snell et al., 2017) and MAML (Finn et al., 2017), and 2) Few-Shot Fine-Tuning (Chen et al., 2019).
 Baselines.
-The selection strategies are compared with 2 baselines: 1) Classic selection, where a new set of samples is randomly selected for each new task, potentially covering the whole datasets when evaluating on enough tasks; and 2) Random selection, where only a single set of 5 samples per class is randomly selected at the start and used for every single single task, representing a more realistic setting with limited budget.
-In-context learning is also compared with the recently proposed LENS (Li & Qiu, 2023) strategy that is in-context learning specific.
-Each selection strategy is run over 10 random seeds to determine their sensitivity to the effects of randomness (e.g., randomly initialised seeds).
-In addition, each experiment is repeated 5 times for different data splits and an average accuracy over these runs is reported (if not specified otherwise).
-Further experimental setup details are in Appendix D.
-4.1 Impact of Single-Property Selection Strategies
-| Strategy | Image Data - FSL | Text Data - FSL | Text Data - ICL | |||||
-| ProtoNet | MAML | FT | ProtoNet | MAML | FT | Mistral | Zephyr | |
+We compare all selected single-property selection strategies (see also Table 2 in Appendix B) with 2 baselines: 1) Classic selection, where a new set of samples is randomly selected for each new task; and 2) Random selection, where only a single set of 5 samples per class is randomly selected once and used for every single task, representing a more realistic setting with limited budget.
+In-context learning is also compared with the recently proposed state-of-the-art in-context learning specific strategies LENS (Li and Qiu, 2023), Active Prompt (Diao et al., 2024), EXPLORA (Purohit et al., 2024) and CASE (Purohit et al., 2025).
+Each selection strategy is run 10 times with different random seeds to determine their sensitivity to the effects of randomness (e.g., random initialisation or sample order).
+In addition, each experiment is repeated 5 times for different data splits (and the choice of the 200 labelled samples) and an average accuracy over these runs is reported (if not specified otherwise).
+Furthermore, each evaluation run is done on 600 randomly sampled tasks for gradient few-shot learning and 300 tasks for in-context learning (due to high computation costs of LLMs), where each task represents a random subset of 5 classes from the dataset (i.e., to construct the 5-way classification setting, following the standard few-shot learning methodology).
+As such, repeating the evaluation across multiple tasks and seeds deals with any possibility of overfitting and minimises the possible biases and other confounding factor in the evaluation.
+Further experiment setup details are in Appendix C.
+4.1 Impact of Selection Strategies
+In this section, our goal is to answer the following research question: RQ1: What is the impact of single-property selection strategies on the few-shot learning?
+The objective is to analyse the performance increase of the different selection strategies across various datasets and few-shot learning approaches.
+The results are shown in Figure 1 and Table 4 in Appendix D, aggregated over all the datasets (the results for image data are included in Appendix E and for individual datasets in supplementary material).
+We present the results separately for in-context learning and gradient few-shot learning, as we observed a strong dependence of the strategy benefit on the few-shot learning approach used.
+Additional ablation studies for the selection strategies are included in Appendix F.
+The majority of selection strategies fail to outperform the Classic selection in case of gradient few-shot learning.
+In most cases, the strategies lead to performance on par or even lower than the Classic selection (for example diverse samples on text datasets).
+Only specific strategies consistently lead to better performance (e.g., combination of easy to learn and ambiguous samples, or Forgetting on text datasets).
+As such, the random sample selection represents a strong baseline for the gradient few-shot learning.
+Sample selection is more beneficial for in-context learning.
+Many of the selection strategies lead to increase in performance for in-context learning, while only few strategies, such as the popular selection based on similarity or diversity, under-perform the Classic selection.
+At the same time, the increase in performance is higher than with gradient few-shot learning, although with a larger variance in results – average increase of for in-context learning as compared to for gradient few-shot learning.
+In contrast to gradient few-shot learning, the hard to learn samples provide the most benefit for in-context learning.
+The reason may be that hard to learn samples provide a lot of information beneficial for the classification that cannot be easily leveraged in the gradient few-shot learning due to the information from the samples being hard to learn or retain.
+However, as there is no explicit "learning" in the in-context learning use of large language models, such samples can be better leveraged.
+As such, learnability property is a strong indicator of quality for all few-shot learning approaches.
+Selecting samples based on how easy (or hard for in-context learning) it is to learn them or how often they are forgotten consistently leads to the overall highest increase in performance.
+4.2 ACSESS Results and Characteristics
+In this section, we answer the following research question: RQ2: How does the ACSESS method perform in comparison to the in-context learning specific baselines and single-property strategies?
+Besides analysing and comparing the benefit with other strategies, we analyse the different weighting schemes of our method.
+The ACSESS method consistently leads to performance increases across all few-shot learning approaches.
+Our proposed method outperforms all the investigated single-property strategies, leading to an average performance increase of up to percentage points for in-context learning and percentage points for gradient few-shot learning.
+Similar behaviour can be observed for the in-context specific baselines as well (LENS, Active Prompt, EXPLORA, CASE; see Table 4 and Figure 2 for comparison).
+Overall, the ACSESS method is able to outperform these baselines as well, achieving on average performance increase of percentage points over the worst performing LENS method and percentage points over the best performing CASE method.
+In all cases, the increase is statistically significant using the Wilcoxon signed-rank test.
+These results confirm that combining simple and well-established selection strategies leads to performance on par or better than strategies specifically tailored for the task of sample selection for in-context learning.
+Although the in-context learning specific strategies may show their benefit in specific setting (e.g., similarity based selection for translation), they do not generalise as well to other tasks.
+The ACSESS method identifies samples with complementary properties, with higher priority on learnability.
+For the different types of few-shot learning (in-context vs.
+gradient learning), our method identifies and combines different set of selection strategies.
+In case of in-context learning, the combination of hard to learn samples (Cartography) that are least often forgotten (Forgetting) on the decision boundary (Margin) and with highest Entropy provide the most benefit.
+For the gradient few-shot learning approaches on text datasets, the combination of easy and ambiguous samples (Cartography) that are least often forgotten (Forgetting) on the decision boundary (Margin) that provide additional informativeness from the Graph-Cut strategy provides the most benefit.
+Further analysing the identified strategies and their weights, we find the ACSESS method gives higher importance to samples that are learnable, lower importance to informative samples (but with multiple sources for both), while the strategies that focus on representativeness are not included at all.
+This further reinforces the finding that learnability is a strong indicator of quality for few-shot learning.
+The dataset and approach specific weighting leads to the best performance.
+The only exception is the Few-Shot Fine-Tuning approach, where the uniform weighting shows the most benefit.
+At the same time, the weighting strategy that incorporates the random selection often leads to worse performance and higher variance in the results.
+Even though the uniform combination is computationally less expensive than the weighted combination, its performance increase is only slightly lower (average difference of ).
+In case of in-context learning, the uniform weighting is often on par with the best-performing CASE baslines.
+As such, the uniform weighting represents a good trade-off between the performance increase and the computation costs, making it in general a better choice.
+Based on these observations, we can draw practical recommendations.
+If the higher computation cost is not a problem, or one is interested in the highest possible increase in performance, the full ACSESS method should be run – which includes running the full strategy identification for each dataset and model, and using the weighted combination of the strategies.
+The cost can be reduced by aggregating the identification over some of the dimensions, such as across multiple dataset or models, which leads to lower benefit.
+However, if extensive resource are not available, or we want to determine the potential benefit of sample selection, using a uniform combination of the Cartography and Margin (and in some cases Forgetting) selection strategies provides a good approximation.
+This configuration does not incur any additional computation costs, while providing consistent improvement on all the datasets and models, which is only slightly lower than when using the model/dataset specific configuration.
+4.3 Effects of Variable Number of Shots and Dataset Size
+In this section, we answer the following research question: RQ3: How do the number of selected shots and the size of the dataset affect the benefit of sample selection?
+We perform two ablation studies to investigate how the benefits of sample selection strategies change as we: 1) increase or decrease the number of selected shots; and 2) decrease the number of available labelled samples (i.e., the size of the dataset).
+For the first ablation study, we apply the ACSESS method and Random selection to select a different number of shots and compare their benefit with training on all available samples (results aggregated over all datasets are in Figure 3 and for individual datasets in Appendix H).
+For the second ablation study, we apply the ACSESS method and Random selection to select 5 samples per class from dataset subsets of different sizes and compare their benefit to the selection from the full dataset (results aggregated over all datasets are in Figure 4 and for individual datasets in Appendix I).
+The selection strategies provide more benefit on lower number of shots, while their impact with further increase in number of shots is negligible after a certain point.
+On the 1-shot setting, the difference in accuracy between the Random selection and the ACSESS strategy can be as high as or percentage points respectively for the in-context learning and gradient-based few-shot learning.
+In case of in-context learning, the impact is affected by the model used.
+For the older models (Mistral and Zephyr) the highest performance increase, as high as percentage points, is achieved at 20-25 shots.
+Afterwards, we observe a significant drop in performance, achieving similar performance to the 1-shot setting when using 50-shots.
+For newer models (LLaMA, Gemma and Qwen), we observe that increasing number of shots has only a negligible impact after around 5 shots, with the increase in performance around percentage points.
+However, we still observe a drop in performance when selecting 50 or more shots.
+The main reason for this drop-off may be the limited context size of the large language models, making the model degenerate to random guessing as the cross-over is text length dependent.
+For the gradient few-shot learning, the highest performance is achieved at lower number of shots () and then remains more or less constant (with slight decrease in some cases).
+In addition, increasing the number of selected shots in case of gradient-based few-shot learning leads to diminished impact, as the ACSESS strategy regresses to random selection when selecting more than 30-40 shots.
+For in-context learning, we observe this behaviour only after the drop in performance.
+As such, the focus on selecting high quality samples is paramount when selecting only a few labelled samples per class, while the random selection can be safely used with higher number of shots.
+However, labelling more samples and using them as additional shots may not always lead to better performance.
+Sample selection is beneficial even when using dataset with low number of samples.
+Using only 25% (50 samples per class) for in-context learning or 10% of the dataset (20 samples per class) for gradient few-shot learning, sample selection can still identify samples that lead to similar performance as is achieved when using the full dataset.
+Decreasing the size further, the benefit of sample selection starts to decrease as well.
+Using only 10 samples per class, the benefit decreases by approximately (e.g., from performance increase of percentage points to only ), and then quickly deteriorates to the Random selection (as the number of samples to choose from is too low).
+5 Conclusion
+In this paper, we investigate whether the well-established sample selection objectives can rival the strategies tailored specifically to in-context learning.
+To this end, we propose the ACSESS method for the automatic combination of the sample selection objectives to select samples with complementary properties and perform extensive evaluation and comparison of sample selection strategies on in-context learning and other few-shot learning approaches.
+The experimental results show that although the individual sample selection strategies may not always be beneficial, their combination using our proposed method consistently leads to better performance, in many cases exceeding the performance from in-context learning specific strategies.
+Curiously, we observe that the learnability of the samples is more important for few-shot learning than their informativeness or representativeness, which is not often utilised in prior works.
+Based on these results, the uniform combination of Cartography, Forgetting and Margin selection strategies is a good alternative transferable across all settings.
+Analysing the behaviour of the sample selection when changing the number of selected shots and the size of the dataset, we find that the sample selection provides the most benefit with a low number of shots, while with a higher number of shots, all strategies regress to random guessing.
+In addition, after a certain point, increasing the number of shots used for training no longer leads to improvement in the performance of few-shot learning, but may lead to a decrease in performance in some cases, especially for in-context learning.
+Overall, the benefit of selecting a subset of high-quality samples lies in improving the performance and reducing its variance for in-context learning, and in the reduction of computation cost, while providing a performance increase on the imbalanced and noisy datasets for the remaining few-shot learning approaches.
+Finally, even when the datasets contain only a few available samples per class, curating the samples further can often still lead to benefits, regardless of dataset characteristics or approach used.
+Limitations
+To diminish undesirable randomness effects, we report the results as an average over 5 repeats of data splitting, 10 repeated runs of the selection strategies (resulting in 50 overall evaluation runs), with each run evaluated on 600 (or 300 for in-context learning) tasks.
+Such setting limits the variance in results, but may also obscure the benefit of selection strategies to a certain extent.
+On the other hand, although we focus only on a 5-way few-shot learning setting (using only 5 classes per task), we do not expect significant change in result for most of the few-shot learning approaches when using higher number of classes (such as 10-way or 20-way classification).
+The only exception is the in-context learning, where we observed the dependence on input size and context width of the models.
+With higher number of classes, the observed performance of large language models would start to deteriorate to the performance that can be achieved with a low number of shots.
+In this case, using larger model sizes may lead to better utilisation of the samples and higher performance increase from the selection strategies.
+Similarly to other research papers focusing on sample selection and few-shot learning via in-context learning, we assume an availability of large-enough annotated dataset from which we select samples using different strategies – we use up to 200 samples per class for each dataset for the main experiments.
+As such, we also have larger validation or held-out set available for evaluating the selection strategies and hyperparameters for our proposed methods or for the approaches themselves.
+If we considered the true few-shot learning setting, where we have only unlabelled data available, only the active learning strategies (i.e., Entropy, Margin and Least Confidence) could be used, as they work with unlabelled data only, or a different approach for sample selection should be employed, such as using weak-labels to select the best samples, or assign every single class to all the labels and perform the sample selection that way to select both best performing samples as well as the class for them (as it is done by Chang and Jia (2023)).
+We address this (potentially most significant) limitation by running ablation studies to observe the benefit of sample selection (using our proposed method and random selection) when using smaller subsets of data (i.e., selecting samples using only fraction of the dataset) or using different number of shots per class.
+Before the experiments, we perform basic prompt-engineering for the in-context learning approaches and run hyperparameter optimisation for the gradient few-shot learning models.
+We follow the recommendations for good prompt engineering, using the meta-tags defined for each model (which lead to better performance of the models), while also not using overly-simple or overly-complicated prompts.
+The prompts we use are created based on dataset description, the prompts used in related works, our own experience, and the formats recommended for different models (e.g., taking inspiration from Sun et al.
+(2023)).
+As we do not perform any extensive or automatic prompt-engineering, these prompts may lead to lowered performance and in some cases also lower impact of the sample selection for the few-shot learning setting.
+However, running an extensive prompt-engineering would introduce another controllable parameter which could potentially introduce further bias into the results (e.g., different selection strategies showing different impact for different prompt formats).
+Ethical Considerations
+The experiments in this paper work with publicly available datasets, citing the original authors of each dataset (with the exception of the datasets contained in MetaAlbum, where we cite the creators of the aggregated dataset).
+All the datasets are used in accordance to their license and terms and conditions of use.
+We do not work with any personally identifiable information or offensive content and perform no crowdsourcing nor further data annotation.
+In addition, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the advancement of the field of Machine Learning and Few-Shot learning (including large language models and in-context learning), such as the use of large amount of computation resources.
+Finally, we follow the license terms for all the models we use – all models and datasets allow their use as part of research.
+The large language models (Mistral, Zephyr, LLaMA, Gemma and Qwen), which we use for the experiments, may contain biases and may generate potentially offensive or harmful content.
+However, the authors of the models already reduce this potential bias as much as possible when training the models, while at the same time we limit the output to few tokens and do not release any output of these models, which should further reduce the potential bias and negative impact.
+Impact Statement: CO2 Emissions Related to Experiments
+The experiments presented in this paper used larger amount of compute resources in order to evaluate all the sample selection strategies for all the different few-shot learning approaches.
+Overall, all the experiments (including preliminary experiments whose results are not reported in this paper) were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCO2eq/kWh.
+A cumulative of approximately 2500 hours of computation was performed on hardware of type A100 PCIe 40GB (TDP of 250W).
+Total emissions are estimated to be 270 kgCO2eq of which 0 percent were directly offset.
+These estimations were conducted using the MachineLearning Impact calculator.
+Whenever possible, we tried to reduce the compute resources used as much as possible.
+The most compute resources were used by the large language models – Mistral-7B, Zephyr-7B, LLaMA-3.1-8B (Grattafiori et al., 2024), Gemma-3-4B (Team, 2025) and Qwen-2.5-7B, and their further tuning.
+To reduce the computation costs, we have opted for using their smaller version and in cases when the further tuning is required, we fine-tune only the last classification layer (keeping the rest frozen).
+In addition, for gradient few-shot learning approaches, we opt for smaller models that do not require such an extensive computation costs, while still providing relevant results.
+Finally, whenever possible, we reuse the already performed computations for the selection strategies as well as using the compute optimised version of our proposed method (more information included in Appendix A), such as generating the feature representation for the different samples only a single time and re-using it whenever possible (i.e., using pre-trained models without further tuning for these representations).
+Acknowledgments
+This work was partially supported by the projects funded by the European Union under the EU Horizon 2020: TAILOR, GA No.
+952215, by the European Union under the Horizon Europe: vera.ai, GA No.
+101070093, AI-CODE, GA No.
+101135437; and by the Slovak Research and Development Agency: Modermed, GA No.
+APVV-22-0414.
+The presented results were obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of Society, Operational Program Integrated Infrastructure.
+Appendix A ACSESS Method Supplementary Details
+In this appendix, we provide and discuss supplementary details of our proposed strategy.
+This includes detailed steps of the forward selection in Algorithm 2, backward selection in Algorithm 3 and datamodels selection in Algorithm 4.
+In addition, we provide low-level details how the method works, suggestions for using the method in order to consider its generalisability and reducing its model/dataset dependence, the computational costs of the method (in comparison with other baselines) and how we reduce them to provide compute optimised version of our method, and, finally, suggestions for selecting the best configuration of the ACSESS method based on the compute-performance trade-off of these configurations.
+Some parts of this section are based on the observations from our experiments.
+Considerations for the computation costs of the ACSESS method
+Although the proposed ACSESS method may at first seem to require significantly more computation resources than the single-property selection strategies or the in-context learning specific baselines, there are easy to apply considerations and modifications that increase the compute-efficiency and reduce the costs.
+We define this analysis into two separate parts as they introduce different computation costs: 1) using the method in practice after the optimal set of strategies is already identified; and 2) identifying the optimal set of strategies to combine.
+As all of the commonly used single-property strategies perform the selection by training the model for a few epochs (or just running an inference in order to obtain embeddings in case of Similarity and Diversity) and then using the training dynamics, these strategies can be run at the same time, reusing the same trained model.
+As such, when using the proposed ACSESS method in practice, we need to train the model only a single time and can use it across all the combined strategies.
+Therefore the cost in this case is equivalent to the single-property selection strategies (and lower than the in-context learning specific baselines).
+Similar optimisation can also be used when identifying the optimal set of strategies to combine, where the training of the model is the main computation cost.
+Even though the strategies are combined with different ones, they are designed to assign a specific score to each sample, which can then be reused in the combination.
+In other words, the strategies are needed to be run only once to obtain the scores which are then used as strong approximation for the objectives.
+The only cost that remains is the cost of evaluating the set of selected strategies by each of the combination.
+In case of the gradient-based few-shot learning strategies, this cost is often negligible, as we are using only a small number of samples.
+However, in the case of in-context learning, the cost scales with the size of the model used and the number of combinations that are explored, as the inference itself is quite expensive.
+The cost of this evaluation step is also dependent on the number of samples and the model used (i.e., training on 25 samples vs.
+full dataset; inference on smaller vs.
+larger models).
+Due to this, the computation cost for the ACSESS strategy is higher (by the number of required inferences and the cost of single inference) over the single-property selection strategy.
+At the same time, the cost is lower or on par with in-context learning specific baselines, which use multiple inferences or trainings of the model while selecting the samples (even though such comparison is hard to make and is based only on our empirical observations during the experiments).
+Finally, the computation cost can be further reduced by foregoing the model and dataset specific tuning (e.g., using the ACSESS method on multiple datasets at the same time), as described in the following section.
+The specific cost comparison of the ACSESS with the baselines is presented in Table 1.
+The costs are calculated as the average GPU wall-clock time for each strategy averaged across all datasets.
+| Strategy | GPU time |
+| Similarity/Diversity | 0.56 |
+| Active Learning | 1.02-1.11 |
+| Core-Set | 0.97-1.24 |
+| LENS | 6.19 |
+| Active Prompt | 5.37 |
+| EXPLORA | 4.98 |
+| CASE | 2.04 |
+| ACSESSWeighted | 2.05 |
+| ACSESSUniform | 1.48 |
+Considerations for using the ACSESS method and reducing its model/dataset dependence
+Similarly to hyperparameter optimisation, we suggest to identify and combine the different strategies using the validation set to prevent any data leakage (i.e., evaluating the performance of different strategies using the validation set).
+At its core, the method is designed to be run separately for each model and dataset.
+Although such setting leads to the most benefit and highest increase in performance, this makes it significantly model/dataset dependent and requires significantly more computation resources.
+In order to reduce the model/dataset dependence, the different steps of the method be run across multiple models or datasets.
+This may be best illustrated on the identification of strategies and further weighting using the Datamodels method.
+Instead of having a separate set of combinations of strategies for each dataset a model, we can concatenate them together to create a single dataset.
+After training the LASSO method on this concatenated dataset, we obtain the relevant strategies and their weights for all the dataset and models.
+Similar modification can be applied to the forward and backward selection as well, where we evaluate the added/removed strategy on all the models/datasets, aggregate the score and select the strategies on this aggregated score.
+However, as we observed a strong dependence on the few-shot approach and the data modality, we suggest to run the ACSESS method separately for each few-shot learning approach and modality.
+In our experiments we also follow this suggestion and aggregate only over the datasets, as we have noticed only negligible changes when running the ACSESS method separately for each dataset (i.e., in some cases, the set of identified strategies we report in the results included 1 more strategy that performed really well on that specific dataset).
+Running the ACSESS strategy on even higher aggregation (over all the approaches, models and modalities) is not recommended as our preliminary experiments indicate that it would not lead to significant performance benefit and the method could fail to identify any relevant selection strategy in some cases.
+Finally, foregoing the specific model/dataset weighted combination and instead using the uniform combination (or the weights provided by the Datamodels approach across multiple datasets) can increase the transferability and generalisability of the ACSESS method further.
+Choosing the best configuration for the ACSESS method
+Based on the results of our experiments, to get the best performance benefit of the ACSESS method, running the full strategy identification, either separately for each model and dataset or aggregated across datasets, and using the weighted combination of these strategies should be used.
+This configuration provides the best results across all datasets, approaches and models, with the exception of Few-Shot Fine-Tuning.
+However, this may require extensive computation resources.
+If such resources are not available, or if we want to quickly determine the potential benefit of the sample selection, using a uniform combination of the Cartography and Margin (and in some cases Forgetting) selection strategies provides a good approximation.
+This configuration does not incur any additional computation costs, while providing consistent improvement on all the datasets and models, which is only slightly lower than when using the model/dataset specific configuration.
+Although, running the full method should always provide the most benefit, this configuration (uniform combination of the 3 best performing strategies identified in our experiments) is a good alternative that is transferable across all settings.
+Appendix B Sample Selection Strategies Details
+| Group | Strategy | Description | Reference | |
+| Informativeness | Similarity | Similarity | Most similar samples. | — |
+| Similarity | Diversity | Most diverse/dissimilar samples. | — | |
+| Active Learn. | Entropy | Highest entropy over class probabilities. | (Park et al., 2022) | |
+| Active Learn. | Margin | Lowest difference in class probability between the 2 most probable classes. | (Park et al., 2022) | |
+| Active Learn. | Least Confidence | Class with lowest assigned probability. | (Park et al., 2022) | |
+| Active Learn. | Loss | Highest loss. | (Park et al., 2022) | |
+| Core-Set | Contrastive Active Learning (CAL) | Predictive likelihood that diverges most from the neighbourhood as determined by KL divergence. | (Margatina et al., 2021) | |
+| Core-Set | DeepFool | Smallest amount of perturbation to change class. | (Ducoffe and Precioso, 2018) | |
+| Core-Set | GraNd | Highest contribution to decline of training loss. | (Paul et al., 2021) | |
+| Core-Set | Graph-Cut | Submodularity function that measures diversity and informativeness. | (Iyer and Bilmes, 2013; Iyer et al., 2021) | |
+| Representativeness | Core-Set | Herding | Minimising distance of subset centre to full dataset. | (Welling, 2009; Chen et al., 2010) |
+| Core-Set | KCenter | Minimising distance of every sample in subset to full dataset. | (Sener and Savarese, 2018; Agarwal et al., 2020) | |
+| Core-Set | CRAIG | Gradients representative of full dataset. | (Mirzasoleiman et al., 2020) | |
+| Core-Set | Glister | Bi-level optimisation maximising log-likelihood. | (Killamsetty et al., 2021) | |
+| Learnability | Core-Set | Forgetting | How often the samples are forgotten. | (Toneva et al., 2018) |
+| Cartography | Cartography | How easy it is to learn the samples.
+Due to no obvious consensus, all of easy, ambiguous, hard and combination of easy and ambiguous samples are considered.
+| (Swayamdipta et al., 2020; Zhang and Plank, 2021) |
+In this appendix, we provide further details about the different single-property sample selection strategies that we evaluate in this paper, as well as for the baselines.
+In addition, we specify the hyperparameters of these methods and how we use them for selecting the set of samples.
+The high-level overview of the single-property strategies is included in Table 2.
+The idea behind the uncertainty-based and active learning selection strategies is that the most informative samples with higher impact are the ones that the model is least certain about (Coleman et al., 2020; Park et al., 2022).
+These strategies select samples iteratively over multiple steps.
+In our experiments, we select 1 sample for every class in individual step.
+In addition, the first sample for every class is chosen randomly, as such setting was observed to lead to better active learning results (Park et al., 2022).
+When using these methods for combination, the samples that are selected after running all the required steps (e.g., 4 steps when choosing 5-shots) use the score assigned to them at their respective step when they were chosen, while the remaining samples are assigned the average score over all the selection steps.
+The Entropy strategy measures the entropy for each of the samples over the predicted class probabilities (which is used as the sample score) and selects the sample with highest entropy.
+The Margin (Breaking Ties) strategy assigns to each sample the difference between the probability of the two most probable classes.
+The samples with lowest such difference are selected as those are the ones the model is most uncertain about (and which can be considered on the decision boundary of the model).
+The Least Confidence strategy simply selects the samples where the most likely class has the lowest probability assigned to it.
+Finally, the Loss strategy selects samples with the highest loss.
+The idea behind the core-set selection strategies is to select subset of samples that are representative of the whole dataset (i.e., the subset has similar properties to the full set).
+The included strategies select samples based on different criteria, such as error/loss, gradient or submodularity.
+These strategies select the full set of samples at the same time according to the assigned scores.
+As such, when using these strategies for combination, the full sample scores are used from this single iteration.
+The Decision Boundary based methods select samples the are close to the decision boundary, as these represent the samples that are hard to separate and should provide the most information to the models.
+The Contrastive Active Learning (CAL) (Margatina et al., 2021) selects the samples whose predictive likelihood diverges the most from their neighbourhood as determined by the KL divergence.
+The DeepFool strategy (Ducoffe and Precioso, 2018) approximates the samples on boundary by perturbing the samples until the predicted class changes and selects the samples that require the smallest perturbation.
+The idea behind Error or Loss based methods is that the samples that contribute more loss or error to the training of model have higher importance and should contribute more to the training when selected.
+The GraNd strategy (Paul et al., 2021) calculates the average contribution of each sample to the decline of the training loss at early stage of training across several independent training runs (calculating gradient norm expectation of the sample).
+The Forgetting strategy (Toneva et al., 2018) simply calculates how often the specific sample is incorrectly classified after being classified correctly in the previous epoch.
+The number of such forgetting events is used as the sample score and samples that are forgotten the most often (or least often) are selected.
+In our evaluation of single-property strategies, we consider only the setting of most often forgotten samples (as it has shown the best performance in the preliminary experiments).
+However, when using the strategy for combination, we also consider the setting of selecting the least forgotten ones.
+The Submodularity based methods use submodular functions (Iyer and Bilmes, 2013; Iyer et al., 2021), which naturally measure the informativeness and diversity of a subset of samples.
+We evaluate the selection based on Graph-Cut, as the method performed the best from all the different submodular functions in the preliminary experiments.
+The Geometry based methods assume that samples that are close in the feature space have similar properties.
+As such, these methods try to remove samples that are close to each other, as they provide redundant information.
+Removing enough of such samples then leads to a small subset of sample that cover the properties of the whole dataset.
+The Herding strategy (Welling, 2009; Chen et al., 2010) iteratively and greedily adds one sample to the subset which minimises the distance between the centre of the subset and the centre of the whole dataset.
+On the other hand, the KCenter strategy (Sener and Savarese, 2018; Agarwal et al., 2020) selects sample that minimises the largest distance between each sample already in the subset and its closest sample not yet included in subset.
+The Gradient Matching based methods try to find a subset of samples whose gradient can imitate the whole dataset.
+The CRAIG strategy (Mirzasoleiman et al., 2020) converts the problem into maximisation of a monotone submodular function which is then solved greedily.
+Finally, the Bilevel Optimisation based methods convert the problem into a bilevel optimisation, with different objectives in the outer and inner level.
+The Glister strategy (Killamsetty et al., 2021) uses a validation set on the outer optimisation, where the subset selection is used as objective, and log-likelihood in the inner optimisation, which handles model parameter optimisation.
+In addition, we also evaluate and use sample selection based on the Cartography (Swayamdipta et al., 2020; Zhang and Plank, 2021), which measures how easy or hard it is to learn different samples.
+This ease of learning is determined by observing the training dynamics of all the samples across few epochs, looking at the average confidence/probability of the correct class and the variance of this confidence.
+The samples with high confidence and low variance are considered to be the easy to learn samples.
+At the same time the samples with small confidence and small to medium variance are considered the hard to learn ones.
+The remaining samples are considered to be ambiguous (medium confidence or samples with high variance).
+We explore four different settings in our experiments and choose the samples accordingly: 1) Easy samples, where we sort the samples based on confidence and choose the top-K samples with highest confidence for each class; 2) Hard samples, where we sort the samples based on confidence and choose the bottom-K samples for each class (i.e., the lowest confidence samples); 3) Ambiguous, where we first calculate average confidence and standard deviation of all the samples, select the samples whose confidence is close around the average confidence (in the interval defined by 1 standard deviation around the average confidence), and then randomly sample from them; and 4) Easy + Ambiguous, where we choose half of the samples from the easy set and the other half from ambiguous set.
+All the approaches mentioned so far perform the selection using a model that is trained for some epochs.
+The number of epochs the training is done for significantly affects the outcome of different strategies.
+For this reason, we search for an optimal number of epochs by evaluating the selection strategies and the samples they select on a single run.
+The optimal number of epochs determined by this search is 10% of the full epochs used for the few-shot fine-tuning models.
+This number represents a good trade-off between the computational cost of the methods and the quality of samples.
+Running the training for full number of epochs may result in better selection of samples, but we consider such setting to not be representative.
+In addition, we use the same number of epochs for each strategy to keep the comparison as representative as possible.
+In addition, we also search for an optimal set of hyperparameters for the selection strategies (such as the learning rate, the optimiser used or batch size).
+The optimal set of hyperparameters that are used throughout the experiments is the use of the Adam optimiser with Cosine Annealing learning rate scheduler, with learning rate of 0.01 for image data and 0.0001 for text data and a batch size of 64.
+To keep the results as comparable as possible, each selection strategy uses the same base model as the one used in the main experiments, i.e., ResNet-18 (He et al., 2016) pretrained on Imagenet dataset (Deng et al., 2009) for image data and pretrained BERT-base (Devlin et al., 2019) for text data in case of gradient few-shot learning, and the same large language models (Mistral (Jiang et al., 2023) and Zephyr (Tunstall et al., 2023)), but with only their last classification layer being learned, for in-context learning.
+In case the selection strategy requires sample features, we use the output of last fully-connected layer (the input of classification layer) as the feature vector similar to Guo et al.
+(2022).
+We also evaluate the sample selection based on Similarity and Diversity as it represents popular approach for in-context learning.
+To select the samples, we use the cosine similarity between the feature representation (obtained from the penultimate layer) of the samples.
+For in-context learning, we instead the most similar (for Similarity) or the most dissimilar (for Diversity) samples for each test sample as is common for this approach (Sun et al., 2023).
+To select the samples for gradient few-shot learning, we first randomly select one sample for each class.
+Afterwards, we expand the set of selected samples by iteratively adding the most similar (for Similarity) or most dissimilar (for Diversity) sample to the average feature representation of the set of selected samples for each class (i.e., if we already have 5 selected samples in the specific class, we calculate the similarity to the representation of each of these 5 samples and take an average over these similarities as the final score).
+Finally, for the in-context learning specific baselines, when these baselines need a base model for optimising the selection (e.g., evaluating the selected samples as part of the method or running different steps), we use all the same models as during all of the experiments (Mistral, Zephyr, LLaMA, Gemma and Qwen).
+In other words, we select a set of samples for each of the models as done for the gradient-based few-shot learning as well.
+This also includes running a hyperparameter search informed by the recommendations of the authors to identify the optimal setup for the search.
+In addition, whenever necessary, we modify the source code to allow for running the method.
+For the core-set selection strategies, we use the modified implementation of the methods released by Guo et al.
+(2022) and its extension provided by Park et al.
+(2022) for active learning methods.
+For the Cartography strategy, we use the modified implementation released by the original authors Swayamdipta et al.
+(2020).
+Appendix C Experiments Setup: Further Details
+The experiments and evaluation of different sample selection strategies are done on the within-domain few-shot learning, where the classes in training, validation and testing datasets all come from the same domain (e.g., from the same dataset).
+Each evaluation run is done on 600 randomly selected tasks for the gradient few-shot learning approaches and 300 tasks for the in-context learning approaches (due to high computation costs of LLMs).
+Each task is a 5-way 5-shot (or -shot for experiments where we change the number of shots) classification task evaluated on 16 test samples per class, following the standard few-shot learning methodology and the evaluation from Ullah et al.
+(2022).
+The performance of a single evaluation run is reported as the mean accuracy across these tasks and the different splits of data into labelled/unlabelled and train-validation-test sets (we use 5 splits of data, where in each we split the available data into train-validation-test sets and select 200 samples per class for training, as mentioned in the main part of the paper).
+These scores are then used to report the findings in the main experiments, either as mean and standard deviation over multiple repeated runs of the specific sample selection strategy or over the results from different models and datasets to produce the aggregated figures.
+As such, we do not explicitly report the standard deviation caused by the different data splits or different sampled tasks.
+However, the observed deviation from these two factors is quite small (on average up to in terms of accuracy), as it is calculated over large number of repeats.
+This is also the reason why we can observe zero standard deviation on some of the sample selection strategies.
+| Datasets | Prompt Format | ||||
+| News classification |
+| Intent classification |
+| Dataset | Verbaliser | ||||
+| 20 News Group | {IBM, Middle East Politics, Windows XP, Motorcycles, Medicine, For Sale, Religion, MS Windows, Baseball, Auto, Hockey, Mac, Graphics, Christianity, Guns, Electronics, Space, Crypto, Atheism, Politics} | ||||
+| News Category | {Politics, World News, Parenting, Money, Wellness, Business, Weddings, Entertainment, Impact, Black Voices, Queer Voices, Crime, Divorce, Food and Drink, Worldpost, Parents, Travel, The Worldpost, Healthy Living, Taste, Media, Culture and Arts, Style, Weird News, Style and Beauty, Comedy, Home and Living, Sports, Environment, Education, Good News, Fifty, Women, Science, Arts and Culture, US News, Green, Tech, Religion, Latino Voices, College, Arts} | ||||
+| ATIS | {Flight, Abbreviation, City, Airfare, Ground Service, Ground Fare, Airline, Flight Number, Aircraft, Distance, Capacity, Flight Time, Quantity, Airport} | ||||
+| {Road Condition, Departure, Duration, Event, Arrival, Directions, Distance, Traffic, Update, Combine, Route, Location} | |||||
+| HWU-64 | {Set Alarm, Definition, Play Music, Set Calendar, Play Radio, Confirm, Quirky, Send Email, Currency, Like, Social Query, News, Neutral, Stop, Calendar Query, Alarm Query, Recommend Movie, Play Audiobook, Weather Query, Praise, Social Post, Play Podcast, Negate, Recipe, Affirm, Remove Calendar, Recommend Location, Remove List, Explain, Email Query, Maths, Stock, Transport Query, Order Takeaway, Datetime Query, Repeat, Factoid, Taxi, Add List, Add Contact, Recommend Event, Mute Volume, List Query, Ticket, Convert Datetime, Joke, Remove Alarm, Traffic, Volume Up, Takeaway Query, Play Game, Contact Query, Music Query, Volume Other, Volume Down, Music Setting, Greet, Dislike} | ||||
+| SNIPS | {Playlist, Weather, Event, Musing, Creative Work, Rate Book, Book Restaurant} |
+For the in-context learning approaches, we use the instruction-tuned Mistral-7B (Jiang et al., 2023) (instruct v0.1), Zephyr-7B (Tunstall et al., 2023) (alpha version), LLaMA-3.1-8B (Grattafiori et al., 2024), Gemma-3-4B (Team, 2025) and Qwen-2.5-7B (Team, 2024) large language models in full precision (although 4-bit precision has shown minimal impact on evaluation).
+We perform no further training for these models (except in case of sample selection, which is covered in Appendix B).
+Instead we run a basic prompt engineering to find the best performing template for each model.
+In the end, we opted for the single template that performed the best on average for each model and each dataset, which is illustrated in Table 3, along with the verbaliser (i.e., the possible outputs of the model that are mapped to the classes in the dataset).
+Finally, the in-context learning models are set to provide deterministic outputs (controlling the temperature and disabling sampling) and set to generate maximum of 10 tokens.
+In case where multiple words that can be mapped to a dataset class are generated, the output is treated as incorrect as we consider the model to be just hallucinating and not working as intended (this behaviour was mainly observed when using the models in zero-shot or 50-shot setting, where the model either does not understand the task or the context is not sufficient for it to follow the instructions and often either repeats all the classes or generates further sentences).
+For the gradient few-shot learning approaches we use a pretrained BERT-base (Devlin et al., 2019) model for the text data and the ResNet-18 (He et al., 2016) for image data, pretrained on Imagenet dataset (Deng et al., 2009), as base models.
+All the models are further trained on each dataset using the different few-shot learning approaches, namely the Prototypical Networks (Snell et al., 2017), Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) and Few-Shot Fine-Tuning (Chen et al., 2019).
+In case of Few-Shot Fine-Tuning, the further training is done on the concatenation of all the meta-training classes and their corresponding samples, while at the meta-test time, only the last layer is replaced and fine-tuned on the specific task with 5 classes.
+The Prototypical Networks and MAML follow the typical few-shot training setting, training on large number of 5-way few-shot tasks.
+The meta-learning approaches on image data are trained on 15 000 tasks and the Few-Shot Fine-Tuning approach is trained on 15 000 randomly sampled batches of size 16.
+On text data, the number of sampled tasks and batches is increased to 25 000.
+In addition, the MAML approach uses 5 training iteration and 10 testing iterations on image data and 10 training and 15 testing on text data, while Prototypical Networks use only 1 in both cases.
+Finally, all approaches use the Adam optimiser with 0.001 (meta) learning rate and 0.01 base learning rate.
+The full training process is repeated 3 times and the best performing model, selected using a separate validation dataset, is selected for use throughout the reported experiments.
+For the LASSO model we use the alpha value of 0.001.
+All the hyperparameters for all models were determined using a hyperparameter search on each data modality and model using a single validation split.
+For the experiments, we use a modified implementation provided by the MetaAlbum (Ullah et al., 2022), extended to further approaches, models and datasets.
+In addition, we also use the hyperparameters used in the paper as a starting point for the hyperparameter search.
+Appendix D Additional Results: Table Results from Text Datasets
+In this appendix, we provide the results from the main experiments in table form – in Table 4.
+| Strategy | Text Data - ICL | Text Data - FSL | ||||||
+| Mistral | Zephyr | LLaMA | Gemma | Qwen | ProtoNet | MAML | FT | |
+| Classic | ||||||||
 | Random | ||||||||
-| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||
 | Similarity | ||||||||
@@
 | CartographyEasy+Ambig. | ||||||||
+| LENS | N/A | N/A | N/A | |||||
+| Active Prompt | N/A | N/A | N/A | |||||
+| EXPLORA | N/A | N/A | N/A | |||||
+| CASE | N/A | N/A | N/A | |||||
 | ACSESSUniform | ||||||||
@@
 | ACSESSWith Random | ||||||||
-In this section, our goal is to answer the following research question: RQ1: What is the impact of single-property selection strategies on the few-shot learning?
-The objective is to analyse the performance increase of the different selection strategies across various datasets and few-shot learning approaches.
-The results are shown in Figure 1 and Table 2, aggregated over all the datasets (the detailed de-aggregated results are included in Appendix F).
-We present the results separately for image and text datasets, as well as for the gradient few-shot learning and in-context learning, as we observed a strong dependence of the strategy benefit on the modality and few-shot learning approach used.
-Majority of selection strategies fail to outperform the Classic selection in case of gradient few-shot learning.
-In most cases, the strategies lead to performance on par or even lower than the Classic selection (for example hard samples on image datasets or diverse samples on text datasets).
-Only specific strategies consistently lead to better performance (e.g., easy to learn samples on image datasets, or Glister on text datasets).
-As such, the random sample selection represents a strong baseline for the gradient few-shot learning.
-Sample selection is more beneficial for in-context learning.
-Larger number of strategies leads to increase in performance, while only few strategies, such as the popular selection based on similarity or diversity, under-perform the Classic selection.
-At the same time, the increase in performance is higher than with gradient few-shot learning, although with a larger variance in results.
-In contrast to gradient few-shot learning, the hard to learn samples provide the most benefit for in-context learning.
-The reason may be that hard to learn samples provide a lot of information beneficial for the classification, which cannot be easily leveraged in the gradient few-shot learning due to the samples being hard to learn.
-However, as there is no explicit “learning” in the in-context learning use of large language models, such samples can be more easily leveraged.
-Learnability property is a strong indicator of quality for few-shot learning.
-The strategies that choose based on how easy it is to learn the samples (Cartography) or how often they are forgotten (Forgetting) consistently lead to overall highest increase in performance.
-The observed improvement is on average as high as percentage point for gradient few-shot learning and percentage points for in-context learning.
-4.2 ACSESS Results and Characteristics
-In this section, we answer the following research question: RQ2: How does the ACSESS method perform in comparison to the single-property strategies?
-Besides analysing the benefit and comparing with other strategies, we analyse the influence of the identified strategies and of the different weighting schemes.
-The ACSESS method consistently leads to performance increase across all few-shot learning approaches.
-Our proposed method outperforms all the investigated single-property strategies, leading to an average performance increase of up to percentage points for gradient few-shot learning and percentage points for in-context learning.
-Comparing with the in-context specific strategy LENS (Li & Qiu, 2023) (see Figure 2 for comparison), which also outperforms all single-property strategies, we observe the increase of ACSESS is statistically more significant (p-value of using the Wilcoxon signed-rank test).
-The ACSESS method identifies samples with complementary properties, with higher priority on learnability.
-The identified selection strategies that are used in the combination are different for the different data modalities (image vs.
-text) and the type of few-shot learning (gradient vs.
-in-context learning).
-For the image datasets, the combination of easy to learn samples (Cartography) that are located on the decision boundary (Margin) provides the most benefit.
-For the gradient few-shot learning approaches on text datasets, the combination of easy and ambiguous samples (Cartography) that are least often forgotten (Forgetting) on the decision boundary (Margin) that provide additional informativeness from the Graph-Cut strategy provides the most benefit.
-In case of in-context learning, the combination of hard to learn samples (Cartography) that are least often forgotten (Forgetting) on the decision boundary (Margin) and with highest Entropy provide the most benefit.
-Further analysing the identified strategies and their weights, we find the ACSESS method gives higher importance to samples that are learnable, lower importance to informative samples (but with multiple sources for both), while the strategies that focus on representativeness are not included at all.
-The dataset and approach specific weighting leads to the best performance.
-The only exception is the Few-Shot Fine-Tuning approach, where the uniform weighting shows the most benefit.
-At the same time, the weighting strategy that incorporates the random selection often leads to worse performance and higher variance in the results.
-Even though the uniform combination is computationally less expensive than the weighted combination, its performance increase is only slightly lower (average difference of ).
-As such, the uniform weighting represents a good trade-off between the performance increase and the computation costs, making it in general a better choice.
-4.3 Effects of Variable Number of Shots
-In this section, we answer the following research question: RQ3: How does the benefit of sample selection change based on the number of shots?
-Our aim is to investigate how the benefit of sample selection strategies changes as we increase or decrease the number of shots that we select.
-In addition, we investigate whether the few-shot learning approaches themselves benefit from larger number of shots.
-The results of this evaluation are included in Figure 3, comparing the benefit of the ACSESS strategy and Random selection over the main experimental setup (5-shot setting).
-The selection strategies provide more benefit on lower number of shots.
-On the 1-shot setting, the difference in accuracy between the Random strategy and the ACSESS strategy can be as high as , or percentage points respectively for the few-shot learning on image data and text data, and in-context learning on text data.
-Increasing the number of selected shots leads to diminished impact, as the ACSESS strategy regresses to random selection when selecting more than 30-40 shots.
-For more detailed analysis we compare all the investigated strategies in the 5-shot and 10-shot setting (distribution of results is shown in Figure 5 in Appendix E).
-This comparison further reinforces our findings, as we find that the difference to random selection is statistically significantly higher in the 5-shot (p-value of using the Wilcoxon signed-rank test).
-Impact of selection strategies and further increase of the number of shots is negligible after certain point.
-For the gradient few-shot learning, the performance after approximately 50 shots remains more or less constant.
-For the in-context learning, the performance even decreases significantly after 20-25 shots, achieving similar performance to the 1-shot setting when using 50-shots.
-The main reason for this drop-off may be the limited context size of the large language models, making the model degenerate to random guessing as the cross-over is text length dependent.
-As such, the focus on selecting high quality samples is paramount when only few labelled samples are available (e.g., working with limited labelling budget), while the random selection can be safely used with enough labels.
-Furthermore, labelling more samples and using them as additional shots for few-shot learning may not lead to better performance.
-4.4 Additional Insights to Sample Selection Strategies
-In this section, we provide additional insights from ablation studies.
-We are mainly interested how the factors such as dataset and model affect the sample selection (based on dataset specific results from Appendix F) and how sensitive the strategies are to their random components (based on the variance from repeated strategy runs in Figure 4).
-Specific strategies show sensitivity to the effects of random seed.
-Changing the random seed for specific strategies may often lead to different set of selected samples and different performance impact.
-The only exception are some of the core set selection strategies (such as Glister or the Graph-Cut), which consistently lead to the selection of the same set of samples.
+Appendix E Additional Experiments: Performance on Image Data
+In this appendix, we present the setup and results of the main experiments done on image data.
+The datasets are chosen similarly to text datasets, where we focus on datasets with different characteristics, such as class imbalance and noisy datasets.
+Specifically, we use 8 datasets from the Meta-Album (Ullah et al., 2022), specifically LR_AM.DOG and LR_AM.AWA for animal classification, VCL.APL for plane classification, HUM_ACT.ACT_410 for classification of human actions, MNF.TEX_DTD for classification of texture, MCR.PRT and MCR.PNU for classification of microscope images, and PLT.FLW for flower classification.
+The results of the main experiments, where we investigate the impact of different single-property sample selection strategies on few-shot learning, as well as provide the comparison with the proposed ACSESS method, are presented in an aggregated form in Table 5 and Figure 5.
+Similarly to experiments on text data, the results for each individual dataset and model are presented in supplementary material.
+| Strategy | Image Data - FSL | ||
+| ProtoNet | MAML | FT | |
+| Random | |||
+| Similarity | |||
+| Diversity | |||
+| Entropy | |||
+| Margin | |||
+| Least Confidence | |||
+| Loss | |||
+| CAL | |||
+| CRAIG | |||
+| DeepFool | |||
+| Forgetting | |||
+| Glister | |||
+| GraNd | |||
+| Herding | |||
+| KCenter | |||
+| Graph-Cut | |||
+| CartographyEasy | |||
+| CartographyAmbiguous | |||
+| CartographyHard | |||
+| CartographyEasy+Ambig. | |||
+| ACSESSUniform | |||
+| ACSESSWeighted | |||
+| ACSESSWith Random | |||
+Although the results on the image modality are slightly different, the main findings hold true.
+Many selection strategies perform worse than the Classic selection (such as hard to learn samples), while only specific ones lead to consistently better performance (e.g., easy to learn samples or those that are forgotten the least).
+A specific new finding that we can draw is that the overall benefit of sample selection is higher for noisy datasets and limited for high-quality datasets.
+Sample selection leads to a considerable performance increase over the Classic selection, on average around percentage points, for the datasets with many noisy samples (as determined by the Cartography) or for datasets with large class imbalance (e.g., difference between lowest and median number of samples per class in few hundred samples), such as HUM_ACT.ACT_410, MCR.PRT or MCR.PNU.
+At the same time, performance increase is quite limited (on average around percentage points) for the datasets with few noisy samples or no extensive class imbalance, such as the VCL.APL, PLT.FLW or NewsCategory.
+When it comes to the proposed ACSESS strategy, we still observe the priority on learnability properties.
+In the case of the image data, the combination of easy to learn samples (Cartography) that are located on the decision boundary (Margin) provides the most benefit.
+This represents the combination of the lowest number of strategies we have observed so far, which indicates that the sample selection strategies are well-designed for image data, or that the image datasets contain high-quality samples that are easier to identify.
+Finally, when it comes to the impact of increasing the number of shots, or using datasets of different sizes, we still observe similar findings on image data (see Figure 6).
+When changing the number of shots, the highest performance is achieved around shots and then remains the same across all of further increases.
+Similarly, when selecting from datasets of smaller sizes, we can observe the same performance when using as few as samples (as compared to using the full dataset).
+At the same time, we observe that the more sophisticated selection strategies outperform random selection most of the time, while degrading to random selection on a high number of shots.
+Appendix F Additional Experiments: Dataset and Model Dependence of Sample Selection Strategies and their Sensitivity to Random Components
+In this appendix, we provide additional insights from ablation studies.
+We are mainly focus on: 1) how the impact of selection strategies is affected by the used models and datasets, i.e., look at the model and dataset dependence of different selection strategies and configurations of our proposed method (based on dataset specific results); and 2) how sensitive the strategies are to their random components, i.e., look at how the set of selected samples changes when the strategies are run multiple times and how this affects the overall performance (based on the variance from repeated strategy runs in Figure 7).
+Specific strategies show large sensitivity to the effects of random seed.
+Changing the random seed for specific strategies often leads to different set of selected samples and different performance impact.
+The core set selection strategies are less affected by this randomness.
+Specific core set strategies (such as Glister or Graph-Cut) consistently select the same set of samples and show no variance in results.
+Other strategies core set strategies (e.g., Forgetting, DeepFool or Herding) show only small variance.
 On the other hand, strategies with random initial condition, such as the similarity selection or some of the active learning strategies, are more sensitive to the effects of the random seed.
-The ACSESS method shows lower sensitivity to repeated runs (with the exception of weighting with random element) but still may lead to different sets of samples based on the random seed.
-Strong dataset dependence for selection strategies.
-The strategy that performs well on one dataset (such as DeepFool on the HUM_ACT.ACT_410 dataset) may perform quite poorly on other datasets (such as MCR.PRT or MCR.PNU).
-Such strategies cannot be used on any dataset without further evaluation.
-Only few strategies perform consistently on the majority of the datasets, such as Herding showing poor performance on all image datasets, or Cartography showing good performance across all datasets and models.
-We do not observe a significant dataset dependence of the ACSESS method, as it identifies the well-performing strategies and leverages their strengths for the specific dataset.
-Strong model dependence and sensitivity to selected samples.
-The strategy performing well for one model (GraNd on MCR.PNU for MAML, or KCenter on 20 News for Zephyr) may result in a decrease of performance for other models (GraNd for Few-Shot Fine-Tuning, or KCenter for Mistral on the same setting).
-Overall, the selected samples resulting from different strategy or from a repeated run of the same strategy affect the Few-Shot Fine-Tuning less, leading to smaller benefit.
-At the same time, the MAML method and the in-context learning models are more affected by the selected samples, leading to more varied results and often higher performance increase.
-Such sensitivity to sample selection was observed in previous works as well (Pecher et al., 2023; Köksal et al., 2022; Zhang et al., 2022; Li & Qiu, 2023).
+The ACSESS method shows lower sensitivity to repeated runs (with the exception of weighting with random element), but still may lead to different sets of samples based on the random seed (as the standard deviation is higher than in core set strategies, i.e., up to for the ACSESS method as opposed to or for majority of the core set strategies).
+Many selection strategies show strong dataset dependence.
+In many cases, the strategies that perform well on one dataset (such as DeepFool on the HUM_ACT.ACT_410 dataset) may perform quite poorly on other datasets (such as MCR.PRT or MCR.PNU).
+Due to this dependence, such strategies cannot be used on any dataset without first evaluating them how well they perform on it.
+Only few strategies perform consistently on the majority of the datasets, such as Herding showing poor performance on all image datasets, or Cartography showing good performance across all datasets and models (while the most beneficial configuration of this strategy is dependent on data modality).
+Finally, we do not observe a significant dataset dependence of the ACSESS method, as it identifies the well-performing strategies and leverages their strengths for the specific dataset.
+This observation holds through for the weighting schemes as well (i.e., both uniform and weighted combinations show consistent performance across datasets).
+Many selection strategies show strong model dependence and sensitivity to selected samples.
+The strategies performing well for one model (GraNd on MCR.PNU for MAML, or KCenter on 20 News for Zephyr) may result in a decrease of performance for other models (GraNd for Few-Shot Fine-Tuning, or KCenter for Mistral on the same setting).
+We do not observe such strong dependence for our proposed method ACSESS, with the exception for Few-Shot Fine-Tuning, where the uniform combination is always better than the weighted combination.
+Overall, having a different set of selected samples, either as a result from using different strategy or from repeated runs of the same strategy, affects the Few-Shot Fine-Tuning less, often leading to smaller benefit.
+At the same time, the MAML method and the in-context learning models (Mistral and Zephyr) are more affected by the selected samples, leading to more varied results and often higher performance increase.
+Such sensitivity to sample selection was observed in previous works as well (Pecher et al., 2024; Köksal et al., 2022; Zhang et al., 2022; Li and Qiu, 2023).
 Curiously, Prototypical Networks gain quite large performance increase from sample selection, but show low sensitivity to the selected samples and variance from repeated runs of sample selection.
 As such, sample selection is more important for MAML and In-Context Learning and less important for Few-Shot Fine-Tuning.
-5 Conclusion
-In this paper, we investigate the impact of larger number of different sample selection strategies on the success of few-shot learning.
-To leverage strengths of these strategies, we propose the ACSESS method for automatic combination of the sample selection strategies to select samples with complementary properties.
-The experimental results show that although the individual sample selection strategies may not always be beneficial, their combination using our proposed method consistently leads to better performance.
-Curiously, we observe that the learnability of the samples is more important for few-shot learning than their informativeness or representativeness.
-Finally, analysing the behaviour of the sample selection in regards to the number of selected shots, we find that the sample selection provides the most benefit on low number of shots, while at the higher number of shots all strategies regress to random guessing.
-In addition, after certain point, increasing the number of available samples no longer leads to improvement in performance of few-shot learning, but may lead to decrease in performance in some cases, especially for in-context learning.
-Acknowledgements
-This work was partially supported by the projects funded by the European Union under the EU Horizon 2020: TAILOR, GA No.
-952215, by the European Union under the Horizon Europe: vera.ai, GA No.
-101070093, AI-CODE, GA No.
-101135437; and by Modermed, a project funded by the Slovak Research and Development Agency, GA No.
-APVV-22-0414.
-Part of the research results was obtained using the computational resources procured in the national project National competence centre for high performance computing (project code: 311070AKF2) funded by European Regional Development Fund, EU Structural Funds Informatization of Society, Operational Program Integrated Infrastructure.
-Ethical Considerations and Impact Statement
-The experiments in this paper work with publicly available datasets, citing the original authors of each dataset.
-All the datasets are used in accordance to their license and terms and conditions of use.
-We do not work with any personally identifiable information or offensive content and perform no crowdsourcing or further data annotation.
-In addition, we are not aware of any potential ethical harms or negative societal impacts of our work, apart from the ones related to the advancement of the field of Machine Learning and Few-Shot learning.
-Finally, we follow the license terms for all the models we use – all models and datasets allow their use as part of research.
-It is possible the large language models (Mistral and Zephyr) we used contain biases and may generate potentially offensive or harmful content.
-However, the authors of the models already reduce this potential bias as much as possible when training the models, while at the same time we limit the output to few tokens and do not release any output of these models, which should further reduce the potential bias and negative impact.
-Impact Statement: CO2 Emissions Related to Experiments
-The experiments presented in this paper used higher compute resources in order to evaluate all the sample selection strategies for all the different few-shot learning approaches.
-Overall, the experiments were conducted using a private infrastructure, which has a carbon efficiency of 0.432 kgCOeq/kWh.
-A cumulative of 2000 hours of computation was performed on hardware of type A100 PCIe 40GB (TDP of 250W).
-Total emissions are estimated to be 216 kgCOeq of which 0 percent were directly offset.
-These estimations were conducted using the MachineLearning Impact calculator presented in (Lacoste et al., 2019).
-Whenever possible, we tried to reduce the compute resources used as much as possible.
-The most compute resources were used by the large language models – Mistral-7B and Zephyr-7B and their further tuning.
-To reduce the computation costs, we use the 4-bit quantised versions of these models and in cases when the further tuning is required, we fine-tune only the last classification layer (keeping the rest frozen).
-In addition, for gradient few-shot learning approaches, we opt for smaller models that do not require such an extensive computation costs, while still providing relevant results.
-Finally, whenever possible, we reuse the already performed computations for the selection strategies, such as generating the feature representation for the different samples only a single time and re-using it whenever possible (e.g., using pre-trained models without further tuning for these representations).
-Appendix A Limitations
-In the experiments for evaluating the impact of different sample selection strategies, we focused on such settings, where the impact of sample selection may be reduced — despite that we observed significant differences in the achieved performance.
-First of all, in within-domain few-shot learning, the samples are selected from the same domain, thus the difference in impact and role of the individual samples may not be as high.
-We may expect that the impact of sample selection in cross-domain few-shot learning may be higher and lead to even larger performance deviations.
-In addition, we focus on a balanced-shot setting, where we select the same number of shots for each class, leading to lower impact of strategies for different classes.
-These additional settings are out of scope of this paper and may be a subject of future research.
-To diminish undesirable randomness effects, we report the results as an average over 5 repeats of data splitting, 10 repeated runs of the selection strategies (resulting in 50 overall evaluation runs), with each run evaluated on 600 tasks.
-Such setting limits the variance in results, but may also obscure the benefit of selection strategies to a certain extent.
-On the other hand, although we focus only on a 5-way few-shot learning setting (using only 5 classes per task), we do not expect significant change in result for most of the few-shot learning approaches when using higher number of classes (such as 10-way or 20-way classification).
-The only exception is the in-context learning, where we observed the dependence on input size and context width of the models.
-With higher number of classes, the observed performance of large language models would start to deteriorate to the performance that can be achieved with a low number of shots.
-Similarly to other research papers focusing on sample selection and few-shot learning via in-context learning, we assume an availability of large-enough annotated dataset from which we select the sample using different strategies.
-This also includes the use of larger validation or held-out set when selecting strategies.
-If we considered the true few-shot learning setting, only the active learning strategies (i.e., Entropy, Margin and Least Confidence) could be used, as they work with unlabelled data only, or a different approach for sample selection should be employed, such as using weak-labels to select the best samples, or assign every single class to all the labels and perform the sample selection that way to select both best performing samples as well as the class for them (as it is done in (Chang & Jia, 2023)).
-Before the experiments, we run hyperparameter optimisation for the gradient few-shot learning models and perform basic prompt-engineering for the in-context learning approaches.
-We follow the recommendations for good prompt engineering, using the meta-tags defined for each model (which lead to better performance of the models), while also not using overly-simple or overly-complicated prompts.
-The prompts we use are created based on dataset description, the prompts used in related works and the formats recommended for different models (e.g., taking inspiration from (Sun et al., 2023)).
-As we do not perform any extensive or automatic prompt-engineering, these prompts may lead to lowered performance and in some cases also lower impact of the sample selection for the few-shot learning setting.
-However, running an extensive prompt-engineering would introduce another controllable parameter which could potentially introduce further bias into the results (e.g., different selection strategies showing different impact for different prompt formats).
-Finally, we also identify factors that we consider to be sufficiently covered, but where the further extension could lead to more generalisable results (nevertheless, also requiring a considerable amount of additional computational resources), such as including more selection strategies, evaluating on more datasets covering other classification tasks (mainly in regards to text datasets) or using larger number of base models for meta-learning and few-shot fine-tuning, or large language models for in-context learning.
-First of all, although we focus on the major selection strategies, which are general for any few-shot learning approach, the work can be extended to include larger number of in-context learning specific strategies or strategies that do not perform selection based on scoring of the samples.
-In addition, even though we run the experiments on 8 image datasets and 6 text datasets, covering different classification tasks with different number of classes, the work could be potentially extended to even larger number of datasets and classification tasks, mainly in case of text modality where we are limited by the number of classes the datasets contain (to sufficiently explore few-shot learning).
-Finally, although we cover all the main groups of few-shot learning approaches, the work could be extended to cover larger number of meta-learning approaches (where we focus on two of them), base models for meta-learning and few-shot fine-tuning (we focus on one model for image data and one model for text data) and large language models for in-context learning (we focus on two of the most recent and currently best performing large language models Mistral and Zephyr).
-However, we do not expect significantly different results and findings, only an improvement of their generalisability.
-Appendix B Few-Shot Learning Problem Formulation
+Appendix G Additional Experiments: Comparison of Sample Selection Impact Between 5-Shot and 10-Shot Setting
+To better explore the effect of increasing number of selected shots on the sample selection strategies, we repeat the evaluation of all the selection strategies in the 10-shot setting (using the same setup and methodology) and compare the results to 5-shot setting.
+For both 5-shot and 10-shot setting, we provide a distribution of the results as a difference to the Classic selection and illustrate the comparison of these distributions in Figure 8.
+Overall, when moving from 5-shot to 10-shots, we observe an increase in the overall performance of the models for all the selection strategies (with the exception of some bad performing ones) as well as the Classic selection.
+At the same time, the increase in the performance of the Classic selection is higher, leading to lower difference and lower impact of the different evaluated selection strategies.
+This finding holds for all strategies across all models and few-shot learning approaches.
+In addition, we also compare the results using Wilcoxon signed-rank test, between the results of 5-shot and 10-shot setting.
+The resulting p-value () indicates that the different to Classic selection is more significant in case of 5-shot setting.
+All in all, these experiments further reinforce the findings from comparison of ACSESS and Random selection, where we observe that the the sample selection progressively leads to lower impact as we increase the number of shots we select.
+Appendix H Additional Experiments: Impact of Increasing Number of Shots for Individual Datasets
+In this appendix, we provide results from investigating the impact of increasing the number of shots on the benefit of sample selection separately for the individual dataset.
+For each dataset, we increase the number of selected shots, starting from 1-shot up to the full dataset.
+The datasets HUM_ACT.ACT_410 and MCR.PRT contain only up to 50 samples per class when all samples are used and the MNF.TEX_DTD only up to 100.
+All the remaining datasets either contain at maximum 200 samples, or are subsampled to 200 samples per class if necessary (this is especially done for the text datasets, which are often significantly larger).
+For in-context learning, we increase the number of shots only up to 50 samples per class, as increasing the number further resulted in significantly lower performance (mainly due to limited context size of the large language models).
+The results for the image datasets are included in Figure 11, for the text datasets using the gradient few-shot learning in Figure 10 and for text datasets using the in-context learning in Figure 9.
+For the majority of the datasets, we observe that the highest performance of the sample selection using the ACSESS method is achieved on lower number of shots and then either remains the same until the full dataset, or gradually decreases as we use more samples.
+Only for specific datasets (LR_AM.DOG and LR_AM.AWA datasets for all approaches; PLT.FLW and SNIPS for MAML; and MNF.TEX_DTD for Prototypical Networks and MAML) we observe continual increase of performance when using sample selection as we increase the number of shots, with the highest performance being achieved only at the full dataset (or slightly beforehand).
+In addition, on specific datasets, not all models behave the same, for example, on the MNF.TEX_DTD dataset, only the Few-Shot Fine-Tuning approach achieves highest performance on lower number of shots (which then monotonically decrease as we increase the number of used samples), while the MAML and Prototypical Networks meta-learning approaches achieve the highest performance only when using almost the full dataset.
+On the other hand, using the Random selection, we often observe monotonic increase in performance as the number of the selected shots increases, with the highest performance being achieved on the full dataset.
+This behaviour can be observed across all datasets and few-shot learning approaches (with exception of in-context learning).
+As such, this further reinforces our findings and main motivation that the data-centric approach can be beneficial for the few-shot learning.
+Curating a subset of the most informative and high-quality samples can often lead to performance on par or better than the one when using all the samples for training.
+Finally, we also observe that the benefit of sample selection is more significant at lower number of shots.
+The difference in performance between the ACSESS method and the Random selection on the lower number of shots is significantly higher and slowly decreases as we use more shots for training.
+The exception is MAML on the PLT.FLW and SNIPS datasets, or Few-Shot Fine-Tuning on PLT.FLW and MCR.PNU datasets, where the performance of the sample selection is similar to the one using random selection even when using small number of shots.
+After a certain point, the sample selection using the ACSESS method degrades to Random selection.
+As such, the focus on selecting high quality samples is paramount when selecting only few labelled samples per class, while random selection can be safely used when using higher number of shots.
+Appendix I Additional Experiments: Selection From Datasets of Different Sizes
+In this section, we analyse how the benefit of sample selection strategies changes as we decrease the size of the dataset available.
+For this ablation study, we apply the ACSESS method and Random selection to select 5 samples per class from dataset subsets of different sizes and compare their benefit to the selection from the full dataset.
+Each dataset is subsampled into multiple sizes, starting with 5 samples per class up to the full dataset.
+Each subsampling is repeated 10 times and an average result over these repeats are reported.
+The full-dataset setting represents the selection from 200 samples per class for each datasets (for datasets with a lower number of samples per class, such as HUM_ACT.ACT_410, MCR.PRT or MNF.TEX_DTD, we observe a constant line after the full dataset was achieved).
+In all cases we perform selection of 5 shots per class using the ACSESS method and Random selection.
+Results aggregated over all datasets are in Figure 4 and for individual datasets for the image datasets are included in Figure 14, for the text datasets using the gradient few-shot learning in Figure 13 and for text datasets using the in-context learning in Figure 12.
+Sample selection is beneficial even when using dataset with low number of samples.
+Using only 25% (50 samples per class) for in-context learning or 10% of the dataset (20 samples per class) for gradient few-shot learning, sample selection can still identify samples that lead to similar performance as is achieved when using the full dataset.
+Decreasing the size further, the benefit of sample selection starts to decrease as well.
+Using only 10 samples per class, the benefit decreases by approximately (e.g., from performance increase of percentage points to only ), and then quickly deteriorates to the Random selection (as the number of samples to choose from is too low).
+For majority of the combinations of datasets and approaches, we can observe that sample selection can achieve similar performance using only a fraction of the dataset.
+For the gradient few-shot learning, this performance is achieved using only 10% of the dataset (20 sample per class), while for in-context learning, this performance is achieved using 25% of the dataset (50 samples per class).
+However, we observe that this strongly depends on the dataset and the model/approach used.
+For example on the ATIS dataset and in-context learning, using only 10 samples per class is enough to achieve the same performance as when selecting from the full dataset, while on the SNIPS dataset the number of samples per class increase up to 100.
+Similarly, on the ATIS dataset with gradient few-shot learning, the MAML a Few-Shot Fine-Tuning approaches achieve the similar performance using only 15-20 samples per class, while Prototypical Networks require the full dataset.
+Curiously, in specific cases, we can observe that running sample selection on subset of the dataset can lead to better performance than when performing selection from the full dataset (such as on the News Category dataset and gradient few-shot learning).
+At the same time, we observe no change in impact of random selection from datasets of different sizes.
+In this case, the performance remains more or less constant (and often significantly below selection using the ACSESS method) throughout all the sizes – only difference in performance stems from the random subsampling.
+Finally, using only 10 samples per class, we can observer that sample selection using the ACSESS method can lead to quite big performance improvement over the random selection.
+Therefore we can conclude that the sample selection is beneficial even when using a small fraction of the dataset size as the one used in the main experiments, although it will lead to a lowered benefit of the sample selection and its impact on the overall performance (which is still higher than when using random selection).
+Appendix J Few-Shot Learning Problem Formulation
 In this section, we introduce the definition of the few-shot learning problem and the difference between gradient-based few-shot learning and few-shot learning via in-context learning.
@@
 After enough episodes are run, the model is evaluated (in a meta-testing phase) using the exact same setup.
-On the other hand, no training phase is present in few-shot learning via in-context learning, but a pretrained large language model is used without any additional parameter updates instead.
-For each new task, a task-specific text prompt is formed, by concatenating task-specific instructions, the samples from the support set along with the their labels, and a specific test sample.
-After being given this prompt, the large language model generate a prediction in a form of word (or multiple words).
-These generated words are then mapped to the task labels and a class is assigned.
+On the other hand, no training phase is present in few-shot learning via in-context learning.
+Instead, a pretrained large language model is used without any additional parameter updates.
+For each new task, a task-specific text prompt is formed by concatenating task-specific instructions, the samples from the support set along with their labels, and a specific test sample.
+After being given this prompt, the large language model generates a prediction in the form of a word (or multiple words).
+These generated words are then mapped to the task labels, and a class is assigned.
 The process is repeated for each sample in the query set.
-For example in case of binary classification, the task-specific instruction may be ”Determine sentiment of following sentence”, the concatenated samples from support set may be formatted as ”sentence, sentiment: positive/negative”, with the word ”positive” mapping to class representing positive sentiment and ”negative” to class representing negative sentiment.
+For example, in the case of binary classification, the task-specific instruction may be "Determine sentiment of following sentence", and the concatenated samples from the support set may be formatted as "sentence, sentiment: positive/negative", with the word "positive" mapping to the class representing positive sentiment and "negative" to the class representing negative sentiment.
 As such, the only supervision for the large language model is from the instruction and the concatenation of support samples.
-Therefore, the model is limited by its max input length.
-At the same time, this approach was found to be sensitive to the selected samples in the support set and their order (Zhao et al., 2021; Zhang et al., 2022; Pecher et al., 2023; Chang & Jia, 2023).
-Appendix C Sample Selection Strategies Details
-In this section, we provide further details about the different single-property sample selection strategies that we evaluate in this paper, as well as for the baselines.
-In addition, we specify the hyperparameters of these methods and how we use them for selecting the set of samples.
-The idea behind the uncertainty-based and active learning selection strategies is that the most informative samples with higher impact are the ones that the model is least certain about (Coleman et al., 2020; Park et al., 2022).
-These strategies select samples iteratively over multiple steps.
-In our experiments, we select 1 sample for every class in individual step.
-In addition, the first sample for every class is chosen randomly, as such setting was observed to lead to better active learning results (Park et al., 2022).
-When using these methods for combination, the samples that are selected after running all the required steps (e.g., 4 steps when choosing 5-shots) use the score assigned to them at their respective step when they were chosen, while the remaining samples are assigned the average score over all the selection steps.
-The Entropy strategy measures the entropy for each of the samples over the predicted class probabilities (which is used as the sample score) and selects the sample with highest entropy.
-The Margin (Breaking Ties) strategy assigns to each sample the difference between the probability of the two most probable classes.
-The samples with lowest such difference are selected as those are the ones the model is most uncertain about (and which can be considered on the decision boundary of the model).
-The Least Confidence strategy simply selects the samples where the most likely class has the lowest probability assigned to it.
-Finally, the Loss strategy selects samples with the highest loss.
-The idea behind the core-set selection strategies is to select subset samples that are representative of the whole dataset and have similar properties.
-The included strategies select samples based on different criteria, such as error/loss, gradient or submodularity.
-These strategies select the full set of samples at the same time according to the assigned scores.
-As such, when using these strategies for combination, the full sample scores are used from this single iteration.
-The Decision Boundary based methods select samples the are close to the decision boundary, as these represent the samples that are hard to separate and should provide the most information to the models.
-The Contrastive Active Learning (CAL) (Margatina et al., 2021) selects the samples whose predictive likelihood diverges the most from their neighbourhood as determined by the KL divergence.
-The DeepFool strategy (Ducoffe & Precioso, 2018) approximates the samples on boundary by perturbing the samples until the predicted class changes and selects the samples that require the smallest perturbation.
-The idea behind Error or Loss based methods is that the samples that contribute more loss or error to the training of model have higher importance and should contribute more to the training when selected.
-The GraNd strategy (Paul et al., 2021) calculates the average contribution of each sample to the decline of the training loss at early stage of training across several independent training runs (calculating gradient norm expectation of the sample).
-The Forgetting strategy (Toneva et al., 2018) simply calculates how often the specific sample is incorrectly classified after being classified correctly in the previous epoch.
-The number of such forgetting events is used as the sample score and samples that are forgotten the most often (or least often) are selected.
-In our evaluation of single-property strategies, we consider only the setting of most often forgotten samples.
-However, when using the strategy for combination, we also consider the setting of selecting the least forgotten ones.
-The Submodularity based methods use submodular functions (Iyer & Bilmes, 2013; Iyer et al., 2021), which naturally measure the informativeness and diversity of a subset of samples.
-We evaluate the selection based on Graph-Cut, as the method performed the best from all the different submodular functions.
-The Geometry based methods assumes that samples that are close to the feature space have similar properties.
-As such, these methods try to remove samples that are close to each to as they provide redundant information and lead to small subset that covers the properties of the whole dataset.
-The Herding strategy (Welling, 2009; Chen et al., 2010) iteratively and greedily adds one sample to the subset which minimises the distance between the centre of the subset and the centre of the whole dataset.
-On the other hand, the KCenter strategy (Sener & Savarese, 2018; Agarwal et al., 2020) selects sample that minimises the largest distance between each sample already in the subset and its closest sample not yet included in subset.
-The Gradient Matching based methods try to find a subset of samples whose gradient can imitate the whole dataset.
-The CRAIG strategy (Mirzasoleiman et al., 2020) converts the problem into maximisation of a monotone submodular function which is then solved greedily.
-Finally, the Bilevel Optimisation based methods convert the problem into a bilevel optimisation, with different objectives in the outer and inner level.
-The Glister strategy (Killamsetty et al., 2021) uses a validation set on the outer optimisation, where the subset selection is used as objective, and log-likelihood in the inner optimisation, which handles model parameter optimisation.
-In addition, we also evaluate and use sample selection based on the Cartography (Swayamdipta et al., 2020; Zhang & Plank, 2021), which measures how easy or hard it is to learn different samples.
-This ease of learning is determined by observing the training dynamics of all the samples across few epochs, looking at the average confidence/probability of the correct class and the variance of this confidence.
-The samples with high confidence and low variance are considered to be the easy to learn samples.
-At the same time the samples with small confidence and small to medium variance are considered the hard to learn ones.
-The remaining samples are considered to be ambiguous (medium confidence or samples with high variance).
-We explore four different settings in our experiments and choose the samples accordingly: 1) Easy samples, where we sort the samples based on confidence and choose the top-K samples with highest confidence for each class; 2) Hard samples, where we sort the samples based on confidence and choose the bottom-K samples for each class (i.e., the lowest confidence samples); 3) Ambiguous, where we first calculate average confidence and standard deviation of all the samples, select the samples whose confidence is close around the average confidence (in the interval defined by 1 standard deviation around the average confidence), and then randomly sample from them; and 4) Easy + Ambiguous, where we choose half of the samples from the easy set and the other half from ambiguous set.
-All the approaches mentioned so far perform the selection using a model that is trained for some epochs.
-The number of epochs the training is done for significantly affects the outcome of different strategies.
-For this reason, we search for an optimal number of epochs by evaluating the selected sample on a single run.
-The optimal number of epochs determined by this search is 10% of the full epochs used for the few-shot fine-tuning models.
-This number represents a good trade-off between the computational cost of the methods and the quality of samples.
-Running the training for full number of epochs may result in better selection of samples, but we consider such setting to not be representative.
-In addition, we use the same number of epochs for each strategy to keep the comparison unbiased and representative.
-In addition to searching for optimal number of epochs, we search for an optimal set of hyperparameters for the methods at the same time (such as the learning rate, the optimiser used or batch size).
-We use the Adam optimiser with Cosine Annealing learning rate scheduler, with learning rate of 0.01 for image data and 0.0001 for text data and a batch size of 64.
-To keep the results as comparable and unbiased as possible, each selection strategy uses the same base model as the one used in the main experiments, i.e., ResNet-18 (He et al., 2016) pretrained on Imagenet dataset (Deng et al., 2009) for image data and pretrained BERT-base (Devlin et al., 2019) for text data in case of gradient few-shot learning, and the same large language models (Mistral (Jiang et al., 2023) and Zephyr (Tunstall et al., 2023)), but with only their last classification layer being learned, for in-context learning.
-In case the selection strategy requires sample features, we use the output of last fully-connected layer (the input of classification layer) as the feature vector similar to (Guo et al., 2022).
-We also evaluate the sample selection based on Similarity and Diversity as it represents popular approach for in-context learning.
-To select the sample, we use the cosine similarity between the feature representation (obtained from the penultimate layer) of the samples.
-To select the samples for gradient few-shot learning, we first randomly select one sample for each class.
-Afterwards, we expand the set of selected samples by iteratively adding the most similar (for Similarity) or most dissimilar (for Diversity) sample to the average feature representation of the set of selected samples for each class (i.e., if we already have 5 selected samples in the specific class, we calculate the similarity to the average vector feature representation of these 5 samples).
-For in-context learning, we instead the most similar (for Similarity) or the most dissimilar (for Diversity) samples for each test sample as is common for this approach (Sun et al., 2023).
-Finally, we also use the in-context learning specific LENS method (Li & Qiu, 2023) as a baseline.
-We use the same models (Mistral and Zephyr) for the selection of samples and run a hyperparameter search informed by the recommendations of the authors to an optimal setup for the search.
-In addition, we use the modified version of their source code for running the method.
-For the core-set selection strategies, we use the modified implementation of the methods released by (Guo et al., 2022) and its extension provided by (Park et al., 2022) for active learning methods.
-For the Cartography strategy, we use the modified implementation released by the original authors (Swayamdipta et al., 2020).
-Appendix D Experimental Setup: Further Details
-The experiments and evaluation of different sample selection strategies are done on the within-domain few-shot learning, where the classes in training, validation and testing datasets all come from the same domain (e.g., from the same dataset).
-Each evaluation run is done on 600 randomly selected tasks for the gradient few-shot learning approaches and 300 tasks for the in-context learning approaches (due to high computation costs of LLMs).
-Each tasks is a 5-way 5-shot (or -shot for experiments where we change the number of shots) classification task evaluated on 16 test samples per class, following the standard few-shot learning methodology and the evaluation from (Ullah et al., 2022).
-The performance of a single evaluation run is reported as the mean accuracy across these tasks and the different splits of data (we use 5 splits of data as mentioned in the main part of the paper).
-These scores are then used to report the findings in the main experiments, either as mean and standard deviation over multiple repeated runs of the specific sample selection strategy or over the different models and datasets.
-As such, the standard deviation of different data splits or different tasks is not reported.
-This is also the reason why we can observe zero standard deviation on some of the sample selection strategies.
-For the gradient few-shot learning approaches we use the ResNet-18 (He et al., 2016) for image data, pretrained on Imagenet dataset (Deng et al., 2009), and a pretrained BERT-base (Devlin et al., 2019) model for the text data as base models.
-All the models are further trained on each dataset using the different few-shot learning approaches, namely the Prototypical Networks (Snell et al., 2017), Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) and Few-Shot Fine-Tuning (Chen et al., 2019).
-In case of Few-Shot Fine-Tuning, the further training is done on the concatenation of all the meta-training classes and their corresponding samples, while at the meta-test time, only the last layer is replaced and fine-tuned on the specific task with 5 classes.
-The Prototypical Networks and MAML follow the typical few-shot training setting, training on large number of 5-way few-shot tasks.
-The meta-learning approaches on image data are trained on 15 000 tasks and the Few-Shot Fine-Tuning approach is trained on 15 000 randomly sampled batches of size 16.
-On text data, the number of sampled tasks and batches is increased to 25 000.
-In addition, the MAML approaches uses 5 training iteration and 10 testing iterations on image data and 10 training and 15 testing on text data, while Prototypical Networks use only 1 in both cases.
-Finally, all approaches use the Adam optimiser with 0.001 (meta) learning rate and 0.01 base learning rate.
-The full training process is repeated 3 times and the best performing model, selected using a separate validation dataset, is selected for use for evaluation.
-| Datasets | Prompt Format | ||||
-| News classification |
-| Intent classification |
-| Dataset | Verbaliser | ||||
-| 20 News Group |
-{IBM, Middle East Politics, Windows XP, Motorcycles, Medicine, For Sale, Religion, MS Windows, Baseball, Auto, Hockey, Mac, Graphics, Christianity, Guns, Electronics, Space, Crypto, Atheism, Politics} |
-| News Category |
-{Politics, World News, Parenting, Money, Wellness, Business, Weddings, Entertainment, Impact, Black Voices, Queer Voices, Crime, Divorce, Food and Drink, Worldpost, Parents, Travel, The Worldpost, Healthy Living, Taste, Media, Culture and Arts, Style, Weird News, Style and Beauty, Comedy, Home and Living, Sports, Environment, Education, Good News, Fifty, Women, Science, Arts and Culture, US News, Green, Tech, Religion, Latino Voices, College, Arts} |
-| ATIS |
-{Flight, Abbreviation, City, Airfare, Ground Service, Ground Fare, Airline, Flight Number, Aircraft, Distance, Capacity, Flight Time, Quantity, Airport} |
-{Road Condition, Departure, Duration, Event, Arrival, Directions, Distance, Traffic, Update, Combine, Route, Location} |
-| HWU-64 |
-{Set Alarm, Definition, Play Music, Set Calendar, Play Radio, Confirm, Quirky, Send Email, Currency, Like, Social Query, News, Neutral, Stop, Calendar Query, Alarm Query, Recommend Movie, Play Audiobook, Weather Query, Praise, Social Post, Play Podcast, Negate, Recipe, Affirm, Remove Calendar, Recommend Location, Remove List, Explain, Email Query, Maths, Stock, Transport Query, Order Takeaway, Datetime Query, Repeat, Factoid, Taxi, Add List, Add Contact, Recommend Event, Mute Volume, List Query, Ticket, Convert Datetime, Joke, Remove Alarm, Traffic, Volume Up, Takeaway Query, Play Game, Contact Query, Music Query, Volume Other, Volume Down, Music Setting, Greet, Dislike} |
-| SNIPS |
-{Playlist, Weather, Event, Musing, Creative Work, Rate Book, Book Restaurant} |
-For the in-context learning approaches, we use instruction tuned Mistral-7B (Jiang et al., 2023) and Zephyr-7B (Tunstall et al., 2023) large language models in the 4-bit quantised mode.
-This significantly reduces the computation costs required, while having low impact on the performance and almost no impact the impact of sample selection.
-We perform no training for these models (except in case of sample selection, which is covered in Appendix C).
-Instead we run a basic prompt engineering to find the best performing template for each model.
-In the end, we opted for the simple template that performed the best on average for each model and each dataset, which is illustrated in Table 3, along with the verbaliser (i.e., the possible outputs of the model that are mapped to the classes in the dataset).
-Finally, the in-context learning models are set to provide deterministic outputs (controlling the temperature and disabling sampling) and set to generate maximum of 10 tokens.
-In case where multiple words that can be mapped to a dataset class are generated, the output is treated as incorrect as the model is just hallucinating and not working as intended (this behaviour was mainly observed when using the models in zero-shot or 50-shot setting, where the model either does not understand the task or the context is not sufficient for it to follow the instructions).
-All the hyperparameters for all models were determined using a hyperparameter search on each data modality and model.
-For the experiments, we use a modified implementation provided by the MetaAlbum (Ullah et al., 2022), extended to further approaches, models and datasets.
-In addition, we also the hyperparameters used in the paper as a starting point for the hyperparameter search.
-Appendix E Comparison of Sample Selection Impact Between 5-Shot and 10-Shot Setting
-To better explore the effect of increasing number of selected shots on the sample selection strategies, we repeat the evaluation of all the selection strategies in the 10-shot setting (using the same setup and methodology) and compare the results to 5-shot setting.
-We provide the comparison of distribution of results, i.e., the difference to Classic selection, in Figure 5.
-Overall, we observe an increase in the overall performance of the models for all the selection strategies (with the exception of some bad performing ones) as well as the Classic selection.
-At the same time, we observe lower difference to the Classic selection for all the evaluated selection strategies.
-This finding holds for all strategies across all models and few-shot learning approaches.
-In addition, we also compare the results using Wilcoxon signed-rank test, between the results of 5-shot and 10-shot setting.
-The resulting p-value () indicates that the different to Classic selection is more significant in case of 5-shot setting.
-All in all, these experiments further reinforce the findings from comparison of ACSESS and random selection, where we observed lower impact of sample selection on larger number of shots.
-Appendix F Dataset Specific Results and Visualisations
-In this appendix, we present the full results for each combination of dataset and model we use in the experiments.
-The results are presented using both figures, showing the difference to Classic selection, as well as tables, showing the overall performance of the model on the specific dataset, the standard deviation from multiple runs of the selection strategy and the difference to the Classic selection (while also showing the results for Classic selection).
-Following is the mapping for different datasets, each presenting results for all the models:
-| LR_AM.DOG | ProtoNet | MAML | Fine-Tuning | |||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||
-| Random | ||||||
-| Similarity | ||||||
-| Diversity | ||||||
-| Entropy | ||||||
-| Margin | ||||||
-| Least Confidence | ||||||
-| Loss | ||||||
-| CAL | ||||||
-| CRAIG | ||||||
-| DeepFool | ||||||
-| Forgetting | ||||||
-| Glister | ||||||
-| GraNd | ||||||
-| Herding | ||||||
-| KCenter | ||||||
-| Graph-Cut | ||||||
-| CartographyEasy | ||||||
-| CartographyAmbiguous | ||||||
-| CartographyHard | ||||||
-| CartographyEasy+Ambig. | ||||||
-| ACSESSUniform | ||||||
-| ACSESSWeighted | ||||||
-| ACSESSWith Random | ||||||
-| LR_AM.AWA | ProtoNet | MAML | Fine-Tuning | |||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||
-| Random | ||||||
-| Similarity | ||||||
-| Diversity | ||||||
-| Entropy | ||||||
-| Margin | ||||||
-| Least Confidence | ||||||
-| Loss | ||||||
-| CAL | ||||||
-| CRAIG | ||||||
-| DeepFool | ||||||
-| Forgetting | ||||||
-| Glister | ||||||
-| GraNd | ||||||
-| Herding | ||||||
-| KCenter | ||||||
-| Graph-Cut | ||||||
-| CartographyEasy | ||||||
-| CartographyAmbiguous | ||||||
-| CartographyHard | ||||||
-| CartographyEasy+Ambig. | ||||||
-| ACSESSUniform | ||||||
-| ACSESSWeighted | ||||||
-| ACSESSWith Random | ||||||
-| VCL.APL | ProtoNet | MAML | Fine-Tuning | |||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||
-| Random | ||||||
-| Similarity | ||||||
-| Diversity | ||||||
-| Entropy | ||||||
-| Margin | ||||||
-| Least Confidence | ||||||
-| Loss | ||||||
-| CAL | ||||||
-| CRAIG | ||||||
-| DeepFool | ||||||
-| Forgetting | ||||||
-| Glister | ||||||
-| GraNd | ||||||
-| Herding | ||||||
-| KCenter | ||||||
-| Graph-Cut | ||||||
-| CartographyEasy | ||||||
-| CartographyAmbiguous | ||||||
-| CartographyHard | ||||||
-| CartographyEasy+Ambig. | ||||||
-| ACSESSUniform | ||||||
-| ACSESSWeighted | ||||||
-| ACSESSWith Random | ||||||
-| HUM_ACT.ACT_410 | ProtoNet | MAML | Fine-Tuning | |||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||
-| Random | ||||||
-| Similarity | ||||||
-| Diversity | ||||||
-| Entropy | ||||||
-| Margin | ||||||
-| Least Confidence | ||||||
-| Loss | ||||||
-| CAL | ||||||
-| CRAIG | ||||||
-| DeepFool | ||||||
-| Forgetting | ||||||
-| Glister | ||||||
-| GraNd | ||||||
-| Herding | ||||||
-| KCenter | ||||||
-| Graph-Cut | ||||||
-| CartographyEasy | ||||||
-| CartographyAmbiguous | ||||||
-| CartographyHard | ||||||
-| CartographyEasy+Ambig. | ||||||
-| ACSESSUniform | ||||||
-| ACSESSWeighted | ||||||
-| ACSESSWith Random | ||||||
-| MNF.TEX_DTD | ProtoNet | MAML | Fine-Tuning | |||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||
-| Random | ||||||
-| Similarity | ||||||
-| Diversity | ||||||
-| Entropy | ||||||
-| Margin | ||||||
-| Least Confidence | ||||||
-| Loss | ||||||
-| CAL | ||||||
-| CRAIG | ||||||
-| DeepFool | ||||||
-| Forgetting | ||||||
-| Glister | ||||||
-| GraNd | ||||||
-| Herding | ||||||
-| KCenter | ||||||
-| Graph-Cut | ||||||
-| CartographyEasy | ||||||
-| CartographyAmbiguous | ||||||
-| CartographyHard | ||||||
-| CartographyEasy+Ambig. | ||||||
-| ACSESSUniform | ||||||
-| ACSESSWeighted | ||||||
-| ACSESSWith Random | ||||||
-| MCR.PRT | ProtoNet | MAML | Fine-Tuning | |||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||
-| Random | ||||||
-| Similarity | ||||||
-| Diversity | ||||||
-| Entropy | ||||||
-| Margin | ||||||
-| Least Confidence | ||||||
-| Loss | ||||||
-| CAL | ||||||
-| CRAIG | ||||||
-| DeepFool | ||||||
-| Forgetting | ||||||
-| Glister | ||||||
-| GraNd | ||||||
-| Herding | ||||||
-| KCenter | ||||||
-| Graph-Cut | ||||||
-| CartographyEasy | ||||||
-| CartographyAmbiguous | ||||||
-| CartographyHard | ||||||
-| CartographyEasy+Ambig. | ||||||
-| ACSESSUniform | ||||||
-| ACSESSWeighted | ||||||
-| ACSESSWith Random | ||||||
-| MCR.PNU | ProtoNet | MAML | Fine-Tuning | |||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||
-| Random | ||||||
-| Similarity | ||||||
-| Diversity | ||||||
-| Entropy | ||||||
-| Margin | ||||||
-| Least Confidence | ||||||
-| Loss | ||||||
-| CAL | ||||||
-| CRAIG | ||||||
-| DeepFool | ||||||
-| Forgetting | ||||||
-| Glister | ||||||
-| GraNd | ||||||
-| Herding | ||||||
-| KCenter | ||||||
-| Graph-Cut | ||||||
-| CartographyEasy | ||||||
-| CartographyAmbiguous | ||||||
-| CartographyHard | ||||||
-| CartographyEasy+Ambig. | ||||||
-| ACSESSUniform | ||||||
-| ACSESSWeighted | ||||||
-| ACSESSWith Random | ||||||
-| PLT.FLW | ProtoNet | MAML | Fine-Tuning | |||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||
-| Random | ||||||
-| Similarity | ||||||
-| Diversity | ||||||
-| Entropy | ||||||
-| Margin | ||||||
-| Least Confidence | ||||||
-| Loss | ||||||
-| CAL | ||||||
-| CRAIG | ||||||
-| DeepFool | ||||||
-| Forgetting | ||||||
-| Glister | ||||||
-| GraNd | ||||||
-| Herding | ||||||
-| KCenter | ||||||
-| Graph-Cut | ||||||
-| CartographyEasy | ||||||
-| CartographyAmbiguous | ||||||
-| CartographyHard | ||||||
-| CartographyEasy+Ambig. | ||||||
-| ACSESSUniform | ||||||
-| ACSESSWeighted | ||||||
-| ACSESSWith Random | ||||||
-| 20 News Group | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||||||
-| Random | ||||||||||
-| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
-| Similarity | ||||||||||
-| Diversity | ||||||||||
-| Entropy | ||||||||||
-| Margin | ||||||||||
-| Least Confidence | ||||||||||
-| Loss | ||||||||||
-| CAL | ||||||||||
-| CRAIG | ||||||||||
-| DeepFool | ||||||||||
-| Forgetting | ||||||||||
-| Glister | ||||||||||
-| GraNd | ||||||||||
-| Herding | ||||||||||
-| KCenter | ||||||||||
-| Graph-Cut | ||||||||||
-| CartographyEasy | ||||||||||
-| CartographyAmbiguous | ||||||||||
-| CartographyHard | ||||||||||
-| CartographyEasy+Ambig. | ||||||||||
-| ACSESSUniform | ||||||||||
-| ACSESSWeighted | ||||||||||
-| ACSESSWith Random | ||||||||||
-| News Category | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||||||
-| Random | ||||||||||
-| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
-| Similarity | ||||||||||
-| Diversity | ||||||||||
-| Entropy | ||||||||||
-| Margin | ||||||||||
-| Least Confidence | ||||||||||
-| Loss | ||||||||||
-| CAL | ||||||||||
-| CRAIG | ||||||||||
-| DeepFool | ||||||||||
-| Forgetting | ||||||||||
-| Glister | ||||||||||
-| GraNd | ||||||||||
-| Herding | ||||||||||
-| KCenter | ||||||||||
-| Graph-Cut | ||||||||||
-| CartographyEasy | ||||||||||
-| CartographyAmbiguous | ||||||||||
-| CartographyHard | ||||||||||
-| CartographyEasy+Ambig. | ||||||||||
-| ACSESSUniform | ||||||||||
-| ACSESSWeighted | ||||||||||
-| ACSESSWith Random | ||||||||||
-| ATIS | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||||||
-| Random | ||||||||||
-| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
-| Similarity | ||||||||||
-| Diversity | ||||||||||
-| Entropy | ||||||||||
-| Margin | ||||||||||
-| Least Confidence | ||||||||||
-| Loss | ||||||||||
-| CAL | ||||||||||
-| CRAIG | ||||||||||
-| DeepFool | ||||||||||
-| Forgetting | ||||||||||
-| Glister | ||||||||||
-| GraNd | ||||||||||
-| Herding | ||||||||||
-| KCenter | ||||||||||
-| Graph-Cut | ||||||||||
-| CartographyEasy | ||||||||||
-| CartographyAmbiguous | ||||||||||
-| CartographyHard | ||||||||||
-| CartographyEasy+Ambig. | ||||||||||
-| ACSESSUniform | ||||||||||
-| ACSESSWeighted | ||||||||||
-| ACSESSWith Random | ||||||||||
-| ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | ||||||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||||||
-| Random | ||||||||||
-| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
-| Similarity | ||||||||||
-| Diversity | ||||||||||
-| Entropy | ||||||||||
-| Margin | ||||||||||
-| Least Confidence | ||||||||||
-| Loss | ||||||||||
-| CAL | ||||||||||
-| CRAIG | ||||||||||
-| DeepFool | ||||||||||
-| Forgetting | ||||||||||
-| Glister | ||||||||||
-| GraNd | ||||||||||
-| Herding | ||||||||||
-| KCenter | ||||||||||
-| Graph-Cut | ||||||||||
-| CartographyEasy | ||||||||||
-| CartographyAmbiguous | ||||||||||
-| CartographyHard | ||||||||||
-| CartographyEasy+Ambig. | ||||||||||
-| ACSESSUniform | ||||||||||
-| ACSESSWeighted | ||||||||||
-| ACSESSWith Random | ||||||||||
-| HWU-64 | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||||||
-| Random | ||||||||||
-| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
-| Similarity | ||||||||||
-| Diversity | ||||||||||
-| Entropy | ||||||||||
-| Margin | ||||||||||
-| Least Confidence | ||||||||||
-| Loss | ||||||||||
-| CAL | ||||||||||
-| CRAIG | ||||||||||
-| DeepFool | ||||||||||
-| Forgetting | ||||||||||
-| Glister | ||||||||||
-| GraNd | ||||||||||
-| Herding | ||||||||||
-| KCenter | ||||||||||
-| Graph-Cut | ||||||||||
-| CartographyEasy | ||||||||||
-| CartographyAmbiguous | ||||||||||
-| CartographyHard | ||||||||||
-| CartographyEasy+Ambig. | ||||||||||
-| ACSESSUniform | ||||||||||
-| ACSESSWeighted | ||||||||||
-| ACSESSWith Random | ||||||||||
-| SNIPS | ProtoNet | MAML | Fine-Tuning | Mistral | Zephyr | |||||
-| Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | Acc. + Std | Diff. | |
-| Classic | ||||||||||
-| Random | ||||||||||
-| LENS | N/A | N/A | N/A | N/A | N/A | N/A | ||||
-| Similarity | ||||||||||
-| Diversity | ||||||||||
-| Entropy | ||||||||||
-| Margin | ||||||||||
-| Least Confidence | ||||||||||
-| Loss | ||||||||||
-| CAL | ||||||||||
-| CRAIG | ||||||||||
-| DeepFool | ||||||||||
-| Forgetting | ||||||||||
-| Glister | ||||||||||
-| GraNd | ||||||||||
-| Herding | ||||||||||
-| KCenter | ||||||||||
-| Graph-Cut | ||||||||||
-| CartographyEasy | ||||||||||
-| CartographyAmbiguous | ||||||||||
-| CartographyHard | ||||||||||
-| CartographyEasy+Ambig. | ||||||||||
-| ACSESSUniform | ||||||||||
-| ACSESSWeighted | ||||||||||
-| ACSESSWith Random | ||||||||||
+Therefore, the model is limited by its maximum input length.
+At the same time, this approach was found to be sensitive to the selected samples in the support set and their order (Zhang et al., 2022; Pecher et al., 2024; Chang and Jia, 2023).